### v2

# The v-numbers and linear presentations of ideals of covers of graphs

Abstract.
Let $G$ be a graph and let $J=I_{c}(G)$ be its ideal of covers. The aims of this work are to study the v-number ${\rm v}(J)$ of $J$ and to study when $J$ is linearly presented using combinatorics and commutative algebra. We classify when ${\rm v}(J)$ attains its minimum and maximum possible values in terms of the vertex covers of the graph that satisfy the exchange property. If the cover ideal of a graph has a linear presentation, we express its v-number in terms of the covering number of the graph. We show necessary and sufficient conditions for the graph $\mathcal{G}_{J}$ of $J$ to be connected. Then, for unmixed König graphs, we classify when $J$ is linearly presented using graph theory, and show some results on Cohen–Macaulay König graphs. If $G$ is unmixed, it is shown that the columns of the linear syzygy matrix of $J$ are linearly independent if and only if $\mathcal{G}_{J}$ has no strong $3$-cycles. One of our main theorems shows that if $G$ is unmixed and has no induced 4-cycles, then $J$ is linearly presented. For unmixed graphs without $3$- and $5$-cycles, we classify combinatorially when $J$ is linearly presented.
Key words and phrases:
v-number, ideal of covers, edge ideals, regularity, linear presentation, König graphs
2010 Mathematics Subject Classification:
Primary 13F20; Secondary 05C22, 05E40, 13H10.
The second and third authors were supported by SNII, México. The first author was supported by a scholarship from CONAHCYT, México.
1. Introduction
Let $G$ be a simple graph with vertex set $V(G)=\{t_{1},\ldots,t_{s}\}$ and edge set $E(G)$. A subset $C\subset V(G)$ is a vertex cover if every edge of $G$ is incident with at least one vertex in $C$. A minimal vertex cover is a vertex cover that is not a proper subset of any other vertex cover. The dual concept of a vertex cover is a stable set, i.e., a subset of $V(G)$ is a stable vertex set if and only if its complement with respect to $V(G)$ is a vertex cover.
Let $S=K[t_{1},\ldots,t_{s}]=\bigoplus_{d=0}^{\infty}S_{d}$ be a polynomial ring over a field $K$ with the standard grading. The monomials of $S$ are denoted $t^{a}:=t_{1}^{a_{1}}\cdots t_{s}^{a_{s}}$, $a=(a_{1},\dots,a_{s})\in\mathbb{N}^{s}$. Let $C_{1},\ldots,C_{r}$ be the minimal vertex covers of $G$ and let $t_{c_{1}},\ldots,t_{c_{r}}$ be their corresponding monomials, that is, $t_{c_{j}}=\prod_{t_{i}\in C_{j}}t_{i}$ for $j=1,\ldots,r$. The following squarefree monomial ideals are associated to $G$.
The edge ideal of $G$, denoted $I(G)$, is the ideal [31]:
$I(G):=(\{t_{i}t_{j}\mid\{t_{i},t_{j}\}\in E(G)\})\subset S,$
and the ideal of covers of $G$, denoted $I_{c}(G)$, is the ideal [12]:
$I_{c}(G):=(t_{c_{1}},\ldots,t_{c_{r}})\subset S.$
The aims of this work are to study the v-number of $I_{c}(G)$ and to study when $I_{c}(G)$ is linearly presented using combinatorics. The contents of this paper are as follows. In Section 2 we present some well-known results that will be needed throughout the paper and introduce some notation.
We say a vertex cover $C$ of $G$ has the exchange property if there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$. Vertex covers with the exchange property always exist, and can be classified using neighbor sets and symmetric differences of vertex covers (Lemma 3.1, Proposition 3.2). The exchange number of $G$, denoted $\alpha_{e}(G)$, is the number of vertices in any smallest vertex cover of $G$ with the exchange property. This number is related to the covering number $\alpha_{0}(G)$ of $G$ which is the number of vertices in any smallest vertex cover of $G$. The v-number of $I_{c}(G)$, denoted ${\rm v}(I_{c}(G))$, is an invariant that was introduced in [6] to study the asymptotic behavior of the minimum distance of projective Reed–Muller-type codes [6, Corollary 4.7]. This invariant is defined for any graded ideal, and can be computed using the software system Macaulay2222 [14] and the algorithm given in [15, Appendix A]. Using a combinatorial description for the v-number of ideals of clutters [17, Theorem 3.5], we obtain:
${\rm v}(I_{c}(G))=\alpha_{e}(G)-1.$
In general, ${\rm v}(I_{c}(G))\geq\alpha_{0}(G)-1$ [17, Lemma 3.16] with equality if and only if there exists a vertex cover $C$ of $G$ with $\alpha_{0}(G)$ vertices satisfying the exchange property (Lemma 3.3).
Letting $J:=I_{c}(G)$, we say $J$ is linearly presented if the monomials $t_{c_{1}},\ldots,t_{c_{r}}$ are of the same degree $\alpha_{0}(G)$ and the kernel of the presentation map $\psi\colon S^{r}\rightarrow S$, $e_{i}^{\prime}\mapsto t_{c_{i}}$, is generated by vectors whose entries are linear forms, where $e_{i}^{\prime}$ is the $i$-th unit vector in $S^{r}$.
Let $G$ be an unmixed graph, that is, $|C_{i}|=\alpha_{0}(G)$ for all $i$. Unmixed graphs are also called well-covered graphs [18]. The graph of $J$, denoted $\mathcal{G}_{J}$, has vertex set $V(\mathcal{G}_{J}):=\{C_{i}\}_{i=1}^{r}$, and the pair $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{J}$ if and only if $|C_{i}\cup C_{j}|=|C_{i}|+1$. According to [3, Proposition 2.1], [8], the ideal $J$ is linearly presented if and only if for all $C_{i},C_{j}\in V(\mathcal{G}_{J})$, $i\neq j$, there is a path in $\mathcal{G}_{J}$ connecting $C_{i}$ and $C_{j}$ whose vertices are contained in $C_{i}\cup C_{j}$ (cf. Proposition 3.4). If $J$ is linearly presented, then $\mathcal{G}_{J}$ is connected but the converse is not true (Example  6.7).
Our first result relates the edges of the graph $\mathcal{G}_{J}$ of $J$ with those of the graph $G$, classifies when ${\rm v}(J)$ attains its minimum possible value in terms of $\mathcal{G}_{J}$, and determine the v-number of $J$ for a wide family of graphs. Recall that a graph without edges is called a discrete graph.
Theorem 3.5. Let $G$ be an unmixed graph and let $J=I_{c}(G)$. The following hold.
  1. (a)
    $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{J}$ if and only if there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C_{i}$, $t_{\ell}\notin C_{i}$ and $(C_{i}\setminus\{t_{k}\})\cup\{t_{\ell}\}=C_{j}$.
  2. (b)
    $\mathcal{G}_{J}$ is not a discrete graph if and only if ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
  3. (c)
    If $J$ is linearly presented, then all minimal vertex covers of $G$ have the exchange property and ${\rm v}(J)=\alpha_{0}(G)-1$.
Given a vertex $t_{i}$ of $G$, we denote the neighborhood of $t_{i}$ by $N_{G}(t_{i})$ and the closed neighborhood of $t_{i}$ by $N_{G}[t_{i}]$. If $N_{G}(t_{k})\subset N_{G}[t_{\ell}]$ for some $\{t_{k},t_{\ell}\}\in E(G)$, then there exists a minimal vertex cover $C$ of $G$ such that $N_{G}(t_{k})\setminus C=\{t_{\ell}\}$ (Proposition 3.6). The converse is not true (Example 6.1). The condition “$N_{G}(t_{k})\subset N_{G}[t_{\ell}]$ for some $\{t_{k},t_{\ell}\}\in E(G)$” first appeared in [26, Proposition 3.1]. This condition implies the inequality ${\rm v}(I_{c}(G))\leq{\rm bight}(I(G))-1$ [26], where ${\rm bight}(I(G))$ is the big height of $I(G)$, that is, the size of a largest minimal vertex cover of $G$.
As usual, the regularity and Krull dimension of the quotient ring $S/I_{c}(G)$ are denoted ${\rm reg}(S/I_{c}(G))$ and $\dim(S/I_{c}(G))$, respectively (Section 2, Definition 2.3). By [23, Theorem 3.31] and [17, Proposition 3.2], one has the inequalities
${\rm bight}(I(G))-1\leq{\rm reg}(S/I_{c}(G))\leq\dim(S/I_{c}(G)),$
with equality on the left if $I(G)$ is sequentially Cohen–Macaulay (cf. Corollary 3.7). By a result of Saha [26, Theorem 3.8] the v-number of $I_{c}(G)$ is a lower bound for the regularity of $S/I_{c}(G)$. Thus, in general, one has the following inequalities
(1.1) $\alpha_{0}(G)-1\leq{\rm v}(I_{c}(G))\leq{\rm reg}(S/I_{c}(G))\leq\dim(S/I_{c}(G))=s-2.$
Let $G$ be a graph and let $\overline{G}$ be its complement. We say $G$ is a codi-graph (complement disconnected graph) if $\overline{G}$ is disconnected. A graph $G$ in which every connected induced subgraph has a disconnected complement is called a cograph. A connected cograph is a codi-graph but the converse is not true (Example 6.3). Let $G_{1},\ldots,G_{k}$ be mutually (vertex) disjoint graphs. Their join, denoted $G_{1}\ast\cdots*G_{k}$, consist of $G_{1}\cup\cdots\cup G_{k}$ and all the lines joining $G_{i}$ with $G_{j}$ for all $i\neq j$. If $k\geq 2$ and $G_{i}$ is a discrete graph for all $i$, we say that $G$ is a complete multipartite graph or a complete $k$-partite graph, and if $|V(G_{i})|=p$ for all $i$ we say that $G$ is homogeneous.
We classify when the v-number of $I_{c}(G)$ attains its maximum possible value in Eq. (1.1) and give a combinatorial description of complete multipartite graphs.
Theorem 3.10. Let $G$ be a graph with $E(G)\neq\emptyset$. The following conditions are equivalent.
  1. (a)
    ${\rm v}(I_{c}(G))=s-2$.
  2. (b)
    ${\rm v}(I_{c}(G))={\rm reg}(S/I_{c}(G))=\dim(S/I_{c}(G))=s-2$ and $G$ is a codi-graph.
  3. (c)
    Every vertex cover of $G$ with the exchange property has exactly $s-1$ vertices.
  4. (d)
    ${\rm v}(I_{c}(H))=|V(H)|-2$ for any induced subgraph $H$ of $G$ with $E(H)\neq\emptyset$.
  5. (e)
    Any induced subgraph $H$ of $G$ with $E(H)\neq\emptyset$ is connected.
  6. (f)
    $G$ is a complete multipartite graph.
We recover the fact that ${\rm v}(I_{c}(G))=s-2$ for complete multipartite graphs [26, Theorem 3.6], and show that $I_{c}(G)$ is Cohen–Macaulay when ${\rm v}(I_{c}(G))=s-2$ (Corollary 3.11). If $G$ is the join of two graphs $G_{1}$ and $G_{2}$, we give a formula to compute the v-number of $G$ in terms of those of $G_{1}$ and $G_{2}$ (Proposition 3.12, Example 6.4). Given an integer $k\geq 1$, let $G_{1}$ be the vertex disjoint union of $k+1$ complete graphs with two vertices, let $G_{2}$ be a discrete graph with one vertex, and let $G=G_{1}*G_{2}$ be their join. Then (Corollary 3.13, cf. [26, Theorem 3.10]):
${\rm reg}(S/I_{c}(G))-{\rm v}(I_{c}(G))=k.$
An edge $e=\{t_{1},t_{2}\}$ of $G$ has the (P) property, if for all edges $\{t_{1},t^{\prime}_{1}\}$ and $\{t_{2},t^{\prime}_{2}\}$ distinct from $e$, one has that $\{t^{\prime}_{1},t^{\prime}_{2}\}$ is an edge of $G$. We characterize when all (resp. none) of the edges of a complete multipartite graph have the (P) property. For these graphs we characterize the unmixed property (Proposition 3.14). If $G$ is a homogeneous complete multipartite graph, it is shown that $\mathcal{G}_{I_{c}(G)}$ is connected if and only if ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$ (Corollary 3.15).
Let $G$ be an unmixed graph and let $r$ be the number of minimal vertex covers of $G$. The linear syzygy matrix of $J=I_{c}(G)$, denoted $\mathcal{LS}(J)$, is an $r\times|E(\mathcal{G}_{J})|$ matrix whose entries are either 00 or have the form ±$\pm t_{i}$, $t_{i}\in S$, and whose columns correspond to linear syzygies of $J$ (Section 2). A $k\times k$ minor of a matrix $\varphi$ with entries in $S$ is the determinant of a square submatrix of $\varphi$ of order $k$. The ideal generated by the $k\times k$ minors of $\varphi$ is denoted by $I_{k}(\varphi)$, the (determinantal) rank of $\varphi$ and height of $I_{k}(\varphi)$ are denoted ${\rm rank}(\varphi)$ and ${\rm ht}(I_{k}(\varphi))$, respectively [25].
If $G$ is an unmixed graph, then $J$ is linearly presented if and only ${\rm rank}(\mathcal{L}\mathcal{S}(J))=r-1$ and ${\rm ht}(I_{r-1}(\mathcal{L}\mathcal{S}(J))\geq 2$ (Proposition 2.12). Any nonzero minor of $\mathcal{LS}(J)$ is of the form ±$\pm t^{a}$, $t^{a}\in S$ (Lemma 2.13, cf. [27, cf. Lemma 3.5]). For unmixed graphs, we show that $\mathcal{G}_{J}$ is connected if and only if ${\rm rank}(\mathcal{LS}(J))=r-1$ (Proposition 4.1).
Let $G$ be an unmixed graph. We describe the induced paths of length $2$ of $\mathcal{G}_{J}$ (Proposition 4.2) and the paths $\{C_{1},\ldots,C_{n}\}$ of $\mathcal{G}_{J}$ such that $C_{i}\subset C_{1}\cup C_{n}$ for all $i$ (Proposition 4.3). An edge $e=\{t,t^{\prime}\}$ of $G$ is called an exchange edge, if there are minimal vertex covers $C$ and $C^{\prime}$ of $G$ such that $C^{\prime}=(C\setminus\{t\})\cup\{t^{\prime}\}$. If $G$ has no isolated vertices and $\mathcal{G}_{J}$ is connected, then $V(G)$ is covered by the exchange edges (Proposition 4.5). If $\mathcal{G}_{J}$ is a tree or a path, we classify when $J$ has a linear presentation (Corollary 4.7) and in the latter case we give a combinatorial criterion (Proposition 4.8). In these cases the reader is refer to [20, Sections 3 and 4] for additional results on the structure of $J$. If $G$ is an unmixed bipartite graph without isolated vertices, we show that $\mathcal{G}_{J}$ is bipartite (Proposition 4.9).
A $3$-cycle $\{C_{1},C_{2},C_{3},C_{1}\}$ of $\mathcal{G}_{J}$ is strong if $C_{2}\setminus C_{1}=C_{3}\setminus C_{1}$. For unmixed graphs, we show that the set $F$ of column vectors of $\mathcal{L}\mathcal{S}(J)$ is a minimal system of generators for $S(F)$, the $S$-module generated by $F$, if and only if $\mathcal{G}_{J}$ has no strong $3$-cycles (Proposition 4.10).
Two distinct vertices $t$ and $t^{\prime}$ of a graph $G$ are called duplicated if $N_{G}(t)=N_{G}(t^{\prime})$. If $\mathcal{G}_{J}$ is connected, we show that $G$ has no duplicated vertices and that every $4$-cycle of $G$ has an edge that does not satisfy the (P) property (Theorem 5.4). The converse is not true (Example 6.6).
Let $P$ be a matching of a graph $G$, that is, $P$ is a set of disjoint edges of $G$. If $P$ is a partition of $V(G)$, we call $P$ a perfect matching. We say that $P$ has the (P) property if each edge of $P$ has the (P) property. A graph $G$ is König if $G$ has a matching with $\alpha_{0}(G)$ edges. König graphs include bipartite graphs by a classical result of König (Theorem 2.6), as well as whisker graphs and very well-covered graphs [24, p. 278]. A perfect matching $P$ of $G$ is of König type if $|P|=\alpha_{0}(G)$ [22]. This notion relates to the fact that very well-covered graphs (Definition 2.8) can be classified using perfect matchings [11, 18] (cf. Proposition 2.9).
A graph $G$ is called Cohen–Macaulay if $I(G)$ is Cohen–Macaulay (Definition 2.3). We show the following two results on Cohen–Macaulay König graphs.
Theorem 5.8. Let $G$ be an unmixed König graph without isolated vertices. The following conditions are equivalent.
  1. (a)
    $G$ is Cohen–Macaulay.
  2. (b)
    $I_{c}(G)$ is linearly presented.
  3. (c)
    $\mathcal{G}_{I_{c}(G)}$ is connected.
  4. (d)
    $G$ has no duplicated vertices.
  5. (e)
    Every induced $4$-cycle has an edge without the (P) property.
  6. (f)
    $G$ has a unique perfect matching.
If $G$ is a very well-covered graph, by a result of Crupi, Rinaldo and Terai [7, Theorem 0.2], $G$ is Cohen–Macaulay if and only if $G$ has a unique perfect matching. Part (a) of Theorem 5.10 gives a combinatorial description of the unique perfect matching of a Cohen–Macaulay König graph without isolated vertices. For a bipartite graph $G$, Zaare-Nahandi [34] revisited the Cohen–Macaulayness of $G$ and proved, by different methods than those of [7], that $G$ is Cohen–Macaulay if and only if $G$ is unmixed and has a unique perfect matching.
Theorem 5.10. If $G$ is a Cohen–Macaulay König graph without isolated vertices, then
  1. (a)
    $G$ has a unique perfect matching $P$ which is the set of exchange edges of $G$, and
  2. (b)
    $\mathcal{G}_{I_{c}(G)}$ is a bipartite graph.
Let $C_{1}\triangle C_{2}$ denote the symmetric difference of $C_{1}$ and $C_{2}$. If $G$ is unmixed and $I_{c}(G)$ is not linearly presented, we show that $G$ has distinct minimal vertex covers $C_{1}$ and $C_{2}$ such that the induced subgraph $G[C_{1}\triangle C_{2}]$ of $G$, with vertex set $C_{1}\triangle C_{2}$, is an unmixed bipartite graph without isolated vertices which is not Cohen–Macaulay (Theorem 5.12). In general the converse is not true (Example 6.8).
We come to one of our main results.
Theorem 5.13. If $G$ is unmixed without induced $4$-cycles, then $I_{c}(G)$ is linearly presented.
As a consequence, for any unmixed graph $G$ without induced $4$-cycles the v-number of $I_{c}(G)$ is equal to $\alpha_{0}(G)-1$ (Corollary 5.14). The converse of Theorem 5.13 is not true (Example  6.5).
The set of all minimal vertex covers of a graph $G$ is denoted by $E(G^{\vee})$. The reason for this notation is explained in Section 2 using the notion of blocker coming from the theory of clutters. Let $G$ be an unmixed graph such that $I_{c}(G)$ is linearly presented. If $G[C_{1}\triangle C_{2}]$ is unmixed and not Cohen-Macaulay, where $C_{1}$ and $C_{2}$ are distinct minimal vertex covers of $G$, we show that there exists $C\in E(G^{\vee})$ such that $C_{1}\cap C_{2}\not\subset C\subset C_{1}\cup C_{2}$ (Theorem 5.15).
We come to another of our main results.
Theorem 5.17. Let $G$ be an unmixed graph without $3$- and $5$-cycles. The following condition are equivalent.
  1. (a)
    $I_{c}(G)$ is linearly presented.
  2. (b)
    If $G[C_{1}\triangle C_{2}]$ is unmixed, where $C_{1},C_{2}\in E(G^{\vee})$, $C_{1}\neq C_{2}$, then it is Cohen–Macaulay.
Consider the following families of graphs:
  1. (a)
    $U$ is the family of unmixed graphs.
  2. (b)
[…]
  3. (c)
    $U_{2}=\{G\in U\mid G\mbox{ has no duplicated vertices}\}$.
  4. (d)
    $U_{3}=\{G\in U\mid\mathcal{G}_{I_{c}(G)}\mbox{ is connected}\}$.
  5. (e)
    $U_{4}=\{G\in U\mid I_{c}(G)\mbox{ is linearly presented}\}$.
  6. (f)
[…]
We show the inclusions $U_{5}\subset U_{4}\subset U_{3}\subset U_{2}\subset U_{1}\subset U,$ and give examples to prove that all inclusions are strict (Proposition 5.18).
In Section 6, we present examples related to some of our results. Then, in Appendix A, we give the procedures for Macaulay2222 [14] that we use to compute the graph of ideals of covers, and regularities, v-numbers, linear syzygy matrices, and ideals of minors of monomial ideals.
For all unexplained terminology and additional information we refer to [16, 23, 30, 33] (for graph theory and edge ideals), and [10, 21, 25] (for commutative rings and ideals of minors).
2. Preliminaries
To avoid repetitions, we continue to employ the notations and definitions used in Section 1. Let $S=K[t_{1},\ldots,t_{s}]=\bigoplus_{d=0}^{\infty}S_{d}$ be a polynomial ring over a field $K$ with the standard grading and let $\mathcal{C}$ be a clutter with vertex set $V(\mathcal{C})=\{t_{1},\ldots,t_{s}\}$, that is, $\mathcal{C}$ is a family of subsets of $V(\mathcal{C})$, called edges, none of which is included in another. Any squarefree monomial ideal is the edge ideal of a clutter [33, pp. 220–221]. An isolated vertex of $\mathcal{C}$ is a vertex that is not in any edge of $\mathcal{C}$. The set of edges of $\mathcal{C}$ is denoted by $E(\mathcal{C})$. The primer example of a clutter is a simple graph $G$. The edge ideal of $\mathcal{C}$, denoted $I(\mathcal{C})$, is the ideal of $S$ generated by all squarefree monomials $t_{e}:=\prod_{t_{i}\in e}t_{i}$ such that $e\in E(\mathcal{C})$. In what follows $I$ denotes the edge ideal of $\mathcal{C}$.
A prime ideal $\mathfrak{p}$ of $S$ is an associated prime of $I$ if $(I\colon f)=\mathfrak{p}$ for some $f\in S_{d}$, where $(I\colon f)$ is the set of all $g\in S$ such that $gf\in I$. The set of associated primes of $I$ is denoted by ${\rm Ass}(I)$. The v-number of $I$, denoted ${\rm v}(I)$, is the following invariant of $I$ [6, Corollary 4.7]:
${\rm v}(I):=\min\{d\geq 0\mid\exists\,f\in S_{d}\mbox{ and }\mathfrak{p}\in{\rm Ass}(I)\mbox{ with }(I\colon f)=\mathfrak{p}\}.$
One can define the v-number of $I$ locally at each associated prime $\mathfrak{p}$ of $I$:
${\rm v}_{\mathfrak{p}}(I):=\mbox{min}\{d\geq 0\mid\exists\,f\in S_{d}\mbox{ with }(I\colon f)=\mathfrak{p}\}.$
A subset $A$ of $V(\mathcal{C})$ is called independent or stable if $e\not\subset A$ for any $e\in E(\mathcal{C})$. The dual concept of a stable set is a vertex cover, i.e., a subset $C$ of $V(\mathcal{C})$ is a vertex cover if and only if $V(\mathcal{C})\setminus C$ is a stable set. A minimal vertex cover is a vertex cover which is minimal with respect to inclusion. If $A$ is a stable set of $\mathcal{C}$, the neighbor set of $A$, denoted $N_{\mathcal{C}}(A)$, is the set of all vertices $t_{i}$ such that $\{t_{i}\}\cup A$ contains an edge of $\mathcal{C}$. For use below, $\mathcal{A}_{\mathcal{C}}$ denotes the family of all stable sets $A$ of $\mathcal{C}$ whose neighbor set $N_{\mathcal{C}}(A)$ is a minimal vertex cover of $\mathcal{C}$. The following result shows that the v-number of a clutter is a combinatorial invariant.
Theorem 2.1.
[17, pp. 9–10] Let $I$ be the edge ideal of a clutter $\mathcal{C}$. If $I$ is not prime, then,
$\mathrm{v}(I)=\min\{|A|\colon A\in\mathcal{A}_{\mathcal{C}}\}\ \mbox{ and }\ (I\colon t_{A})=(N_{\mathcal{C}}(A))\ \forall\,A\in\mathcal{A}_{\mathcal{C}}.$
The number of vertices in any smallest vertex cover of $\mathcal{C}$, denoted $\alpha_{0}({\mathcal{C}})$, is called the covering number of $\mathcal{C}$. The height ${\rm ht}(I)$ of $I$ is $\alpha_{0}(\mathcal{C})$ and the Krull dimension $\dim(S/I)$ of $S/I$ is $s-{\rm ht}(I)$. A clutter is well-covered if every maximal stable set is a maximum stable set. The clutter of minimal vertex covers of $\mathcal{C}$, denoted $\mathcal{C}^{\vee}$, is called the blocker of $\mathcal{C}$. The edge set $E(\mathcal{C}^{\vee})$ of the blocker of $\mathcal{C}$ is the set of minimal vertex covers of $\mathcal{C}$ and the vertices of $\mathcal{C}^{\vee}$ are the vertices of $\mathcal{C}$. The edge ideal of $\mathcal{C}^{\vee}$, denoted by $I_{c}(\mathcal{C})$, is called the ideal of covers of $\mathcal{C}$.
The deletion of a vertex $t_{i}$, denoted by $\mathcal{C}\setminus t_{i}$, is the clutter obtained from $\mathcal{C}$ by deleting the vertex $t_{i}$ and all edges containing $t_{i}$. Let $A$ be a set of vertices of $\mathcal{C}$. The deletion of $A$, denoted $\mathcal{C}\setminus A$, is obtained by successively deleting the vertices of $A$. The induced clutter on $A$, denoted $\mathcal{C}[A]$, is the clutter on $A$ with edge set consisting of all edges of $\mathcal{C}$ contained in $A$.
A graded ideal is called unmixed if all its associated primes have the same height. A clutter is called unmixed if all its minimal vertex covers have the same number of elements. The next result relates minimal vertex covers and associated primes.
Lemma 2.2.
[33, Lemma 6.3.37] Let $C$ be a set of vertices of a clutter $\mathcal{C}$. Then, $C$ is a minimal vertex cover of $\mathcal{C}$ if and only if the ideal of $S$ generated by $C$ is an associated prime of $I(\mathcal{C})$.
Definition 2.3.
Let $I\subset S$ be a graded ideal and let ${\mathbf{F}}$ be the minimal graded free resolution of $S/I$ as an $S$-module:
[…]
The ideal $I$ has a $d$-linear resolution if all maps $\varphi_{i}$, $i\geq 2$, are defined by matrices whose entries are linear forms and all entries of $\varphi_{1}$ are forms of degree $d$. The projective dimension of $S/I$, denoted ${\rm pd}_{S}(S/I)$, is equal to $g$. The Castelnuovo–Mumford regularity of $S/I$ (regularity of $S/I$ for short) is defined as
${\rm reg}(S/I)=\max\{j-i\mid b_{i,j}\neq 0\}.$
If ${\rm pd}_{S}(S/I)$ is the height of $I$, we say the ring $S/I$ and the ideal $I$ are Cohen–Macaulay.
An element $f\in S$ is called a zero-divisor of $S/I$—as an $S$-module—if there is $\overline{0}\neq\overline{a}\in S/I$ such that $f\overline{a}=\overline{0}$, and $f$ is called regular on $S/I$ otherwise. The depth of $S/I$, denoted ${\rm depth}(S/I)$, is the largest integer $\ell$ so that there is a homogeneous sequence $g_{1},\ldots,g_{\ell}$ in $\mathfrak{m}=(t_{1},\ldots,t_{s})$ with $g_{i}$ not a zero-divisor of $S/(I,g_{1},\ldots,g_{i-1})$ for all $1\leq i\leq\ell$.
The Auslander–Buchsbaum formula [33, p. 226]:
(2.1) ${\rm pd}_{S}(S/I)+{\rm depth}(S/I)=\dim(S)=s$
relates the depth and the projective dimension.
Theorem 2.4.
Let $\mathcal{C}$ be a clutter and let $I_{c}(\mathcal{C})$ be its ideal of covers. The following hold.
  1. (a)
    (Eagon-Reiner [9]) $I(\mathcal{C})$ is Cohen–Macaulay if and only if $I_{c}(\mathcal{C})$ has a linear resolution.
  2. (b)
    (Terai [29]) ${\rm reg}(I(\mathcal{C}))=1+{\rm reg}(S/I(\mathcal{C}))={\rm pd}(S/I_{c}(\mathcal{C}))$.
Proposition 2.5.
Let $\mathcal{C}$ be a clutter, let $I(\mathcal{C})$ be its edge ideal and let $d=\alpha(I(\mathcal{C}))$ be the least degree of a minimal generator of $I(\mathcal{C})$. Then, the edge ideal $I(\mathcal{C})$ has a $d$-linear resolution if and only if ${\rm reg}(S/I(\mathcal{C}))=d-1$.
Proof.
Note that, by duality [33, p. 222], $d=\alpha(I(\mathcal{C}))={\rm ht}(I_{c}(\mathcal{C}))=\alpha_{0}(\mathcal{C}^{\vee})$.
$\Rightarrow$) Assume that $I(\mathcal{C})$ has a $d$-linear resolution. Then, by Theorem 2.4(a), $I_{c}(\mathcal{C})$ is Cohen–Macaulay, that is,
${\rm pd}(S/I_{c}(\mathcal{C}))={\rm ht}(I_{c}(\mathcal{C}))=d.$
Hence, by Theorem 2.4(b), $1+{\rm reg}(S/I(\mathcal{C}))={\rm pd}(S/I_{c}(\mathcal{C}))=d$, and ${\rm reg}(S/I(\mathcal{C}))=d-1$.
$\Leftarrow$) Assume that ${\rm reg}(S/I(\mathcal{C}))=d-1$. Then, by Theorem 2.4(b), one has
${\rm pd}(S/I_{c}(\mathcal{C}))=1+{\rm reg}(S/I(\mathcal{C}))=d={\rm ht}(I_{c}(\mathcal{C})).$
Thus, $I_{c}(\mathcal{C})$ is Cohen–Macaulay and, by Theorem 2.4(a), $I(\mathcal{C})$ has a $d$-linear resolution. ∎
Theorem 2.6.
[16, Theorem 10.2] If $G$ is a bipartite graph and $\alpha_{0}(G)$ is its covering number, then $G$ has a matching with $\alpha_{0}(G)$ edges.
Proposition 2.7.
Let $G$ be a König graph without isolated vertices. The following hold.
  1. (a)
    [5, Proposition 15] $G$ is unmixed if and only if $G$ has a perfect matching $P$ of König type with the (P) property.
  2. (b)
    [5, Proposition 28] $G$ is Cohen–Macaulay if and only if $G$ has a perfect matching of König type $P$ with the (P) property such that there are no $4$-cycles with two edges in $P$.
Definition 2.8.
An unmixed graph $G$ without isolated vertices is called very well-covered if $2\alpha_{0}(G)=|V(G)|$.
Proposition 2.9.
A graph $G$ is very well-covered if and only if $G$ is unmixed and has a perfect matching $P$ of König type.
Proof.
This result follows readily from the fact that very well-covered graphs have a perfect matching [18, Lemma 2.1]. ∎
Proposition 2.10.
[5, Corollary 29] Let $G$ be a Cohen–Macaulay König graph without isolated vertices, then $G$ has at least one free vertex.
In what follows $I$ is a monomial ideal of $S$ minimally generated $\{t^{v_{i}}\}_{i=1}^{q}$. We say $I$ is linearly presented if $\deg(t^{v_{i}})=d$ for all $i$ and the kernel of the presentation map $\psi\colon S^{q}\rightarrow S$, $e_{i}\mapsto t^{v_{i}}$, is generated by vectors whose entries are linear forms, where $e_{i}$ is the $i$-th unit vector in $S^{q}$.
If $I$ is generated in degree $d$, the graph of $I$, denoted $\mathcal{G}_{I}$, has vertex set $V(\mathcal{G}_{I}):=\{t^{v_{i}}\}_{i=1}^{q}$, and $\{u,v\}$ is an edge of $\mathcal{G}_{I}$ if and only if $u\neq v$ and $\deg({\rm lcm}(u,v))=d+1$. For all $u,v$ in $V(\mathcal{G}_{I})$, $u\neq v$, let $\mathcal{G}_{I}^{(u,v)}$ be the induced subgraph of $\mathcal{G}_{I}$ with vertex set
$V(\mathcal{G}_{I}^{(u,v)})=\{t^{v_{i}}\mid t^{v_{i}}\mbox{ divides }{\rm lcm}(u,v)\}.$
Proposition 2.11.
[3, Proposition 2.1, Corollary 2.2] Let $I$ be a monomial ideal of $S$ generated in degree $d$. The following conditions are equivalent.
  1. (a)
    The ideal $I$ is linearly presented.
  2. (b)
    $\mathcal{G}_{I}^{(u,v)}$ is connected for all $u,v\in\{t^{v_{i}}\}_{i=1}^{q}$.
  3. (c)
    For all $u,v\in\{t^{v_{i}}\}_{i=1}^{q}$ there is a path in $\mathcal{G}_{I}^{(u,v)}$ connecting $u$ and $v$ whose vertices are minimal generators of $I$ dividing ${\rm lcm}(u,v)$.
Consider the following basic matrices [27]:
  1. (a)
    The syzygy matrix of $I$, denoted $\mathcal{S}(I)$, is the $q\times\binom{q}{2}$ matrix whose columns are the vectors of the form
    $({{\rm lcm}(t^{v_{i}},t^{v_{j}})}/{t^{v_{i}}})e_{i}-({{\rm lcm}(t^{v_{i}},t^{v_{j}})}/{t^{v_{j}}})e_{j}.$
    Each column of $\mathcal{S}(I)$ defines a syzygy of $I$ because $(t^{v_{1}},\ldots t^{v_{q}})\mathcal{S}(I)=(0)$.
  2. (b)
    The linear syzygy matrix of $I$, denoted $\mathcal{L}\mathcal{S}(I)$, is the $q\times|E(\mathcal{G}_{I})|$ matrix whose columns are the vectors of the form $t_{\ell}e_{i}-t_{k}e_{j}$ such that $t_{\ell}t^{v_{i}}=t_{k}{t^{v_{j}}}$. We set $\mathcal{LS}(I)=(0)$ if $I$ has no linear syzygies. Note that $\mathcal{L}\mathcal{S}(I)$ is a submatrix of $\mathcal{S}(I)$.
  3. (c)
    The numerical linear syzygy matrix, denoted by $\mathcal{N\!LS}(I)$, is the $\{0,\pm 1\}$-matrix obtained from $\mathcal{L}\mathcal{S}(I)$ by making the substitution $t_{i}=1$ for all $i$. This matrix has entries in $\{0,\pm 1\}\subset\mathbb{Z}$ if the base field $K$ has characteristic 00.
  4. (d)
    The oriented graph of $\mathcal{G}_{I}$ obtained from $\mathcal{G}_{I}$ by fixing an orientation for each edge of $\mathcal{G}_{I}$ is denoted by $\mathcal{D}_{I}$.
Proposition 2.12.
Let $I$ be a monomial ideal minimally generated by $t^{v_{1}},\ldots,t^{v_{q}}$ and let $\mathcal{S}(I)$ be the syzygy matrix of $I$. The following hold.
  1. (a)
    [1, Theorem 3.4.1] The kernel of the presentation map $\psi\colon S^{q}\rightarrow S$, $e_{i}\mapsto t^{v_{i}}$, is generated by the columns of $\mathcal{S}(I)$.
  2. (b)
    Let $\varphi$ be a $q\times m$ submatrix of $\mathcal{S}(I)$ whose columns minimally generate a graded $S$-module. Then, ${\rm im}(\varphi)=\ker(\psi)$ if and only if ${\rm rank}(\varphi)=q-1$ and ${\rm ht}(I_{q-1}(\varphi))\geq 2$.
  3. (c)
    If $I$ is generated by monomials of degree $d$, then $I$ is linearly presented if and only ${\rm rank}(\mathcal{L}\mathcal{S}(I))=q-1$ and ${\rm ht}(I_{q-1}(\mathcal{L}\mathcal{S}(I)))\geq 2$.
Proof.
(b) As $\varphi$ is a submatrix of the syzygy matrix of $I$, letting $\psi=(t^{v_{1}},\ldots,t^{v_{q}})$, the sequence
$S^{m}\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}S^{q}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S$
is a free complex of graded $S$-modules. Then, by the Buchsbaum–Eisenbud acyclicity criterion [4, Corollary 1], it is seen that ${\rm im}(\varphi)=\ker(\psi)$ if and only if
$q={\rm rank}(\varphi)+{\rm rank}(\psi)={\rm rank}(\varphi)+1\mbox{ and }\ {\rm ht}(I_{q-1}(\varphi))\geq 2.$
(c) This part follows from (a) and (b). ∎
Lemma 2.13.
[27, cf. Lemma 3.5] Let $\mathcal{S}(I)$ be the syzygy matrix of a monomial ideal $I$. Then, any nonzero minor of $\mathcal{S}(I)$ is a monomial with coefficient ±$\pm 1$, i.e., it has the form ±$\pm t^{a}$.
Proof.
We proceed by induction on the size $n$ of the minor. The case $n=1$ is clear because each entry of $\mathcal{S}(I)$ is either zero or has the form ±$\pm t^{a}$, $a\in\mathbb{N}^{s}\setminus\{0\}$. Assume $n\geq 2$. We may assume that the given nonzero minor is formed by the submatrix $M$ with the first $n$ rows and columns of $\mathcal{S}(I)$. Let $M^{\prime}$ denote the $q\times n$ submatrix of $\mathcal{S}(I)$ with the first $n$ columns. By definition of the syzygy matrix of $I$, any column of the latter has exactly two nonzero entries. It follows that the $q-n$ rows of $M^{\prime}$ complementary to $M$ cannot all be zero as otherwise $M$ would be a matrix of syzygies of the initial $n$ monomials $\{t^{v_{1}},\ldots,t^{v_{n}}\}$ of $I$, which is impossible since $\det(M)\neq 0$ and $(t^{v_{1}},\ldots,t^{v_{n}})M=0$ while the entire syzygy matrix of the ideal generated by these monomials has rank $n-1$ (Proposition 2.12).
Thus, there must be a nonzero entry in some row of $M^{\prime}$ complementary to $M$, say, in the $j$-th column of $M^{\prime}$, with $1\leq j\leq n$. By the construction of $\mathcal{S}(I)$, there is exactly one further nonzero entry on the $j$-th column. This entry must belong to $M$ as otherwise $\det(M)=0$. Also, this entry is again monomial with coefficient ±$\pm 1$. Expanding $\det(M)$ by the $j$-th column yields the product of this monomial by the minor of a suitable $(n-1)\times(n-1)$ submatrix of $M$. By induction, this minor has the required form, hence so does $\det(M)$. ∎
Proposition 2.14.
Let $I$ be a monomial ideal generated in degree $d$ and let $c$ be the number of connected components of the graph $\mathcal{G}_{I}$ of $I$. Then
${\rm rank}(\mathcal{LS}(I))={\rm rank}(\mathcal{N\!LS}(I))=q-c.$
In particular, $\mathcal{G}_{I}$ is connected if and only if ${\rm rank}(\mathcal{LS}(I))=q-1$.
Proof.
By Lemma 2.13, we may assume that $K=\mathbb{Q}$. Let $r_{0}={\rm rank}(\mathcal{LS}(I))$ and let $M$ be a $r_{0}\times r_{0}$ submatrix of $\mathcal{LS}(I)$ with $\det(M)\neq 0$. Then, by Lemma 2.13, $\det(M)=\pm t^{a}$ for some $t^{a}\in S$. Thus, making $t_{i}=1$ for all $i$, we get ${\rm rank}(\mathcal{LS}(I))\leq{\rm rank}(\mathcal{N\!LS}(I))$. The reverse inequality follows by definition of $\mathcal{N\!LS}(I)$. Let $\mathcal{D}_{I}$ be the digraph obtained from $\mathcal{G}_{I}$ by choosing an orientation for the edges of $\mathcal{G}_{I}$. The incidence matrix of $\mathcal{D}_{I}$ is equal to $\mathcal{N\!LS}(I)$ and this is a totally unimodular matrix. Hence, by [13, Theorem 8.3.1], the rank of $\mathcal{N\!LS}(I)$ is $q-c$. ∎
Lemma 2.15.
(Inclusion-exclusion [2, p. 38]) Let $A_{1},\ldots,A_{n}$ be finite subsets of $S$, then
$\left|\bigcup_{i=1}^{n}A_{i}\right|=\sum_{i=1}^{n}|A_{i}|-\sum_{i<j}|A_{i}\cap A_{j}|+\sum_{i<j<k}|A_{i}\cap A_{j}\cap A_{k}|\mp\cdots+(-1)^{n-1}\left|\bigcap_{i=1}^{n}A_{i}\right|.$
3. The v-number of ideals of covers
In what follows to avoid repetitions, we continue to employ the notations and definitions used in Sections 1 and 2.
Lemma 3.1.
Let $G$ be a graph with $E(G)\neq\emptyset$. Then, there exists a vertex cover $C$ of $G$ such that $C$ has the exchange property and $|C|=|V(G)|-1$.
Proof.
Pick an edge $\{t_{k},t_{\ell}\}$ of $G$ and let $A=V(G)\setminus\{t_{k},t_{\ell}\}$. Then, $C=A\cup\{t_{k}\}$ is a vertex cover of $G$ such that $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$ and $|C|=s-1=|V(G)|-1$. ∎
Proposition 3.2.
(Exchange property) Let $C$ be a vertex cover of a graph $G$. The following conditions are equivalent.
  1. (a)
    There exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $N_{G}(t_{k})\setminus C=\{t_{\ell}\}$.
  2. (b)
    There exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$.
  3. (c)
    There exist $\{t_{k},t_{\ell}\}\in E(G)$ and a vertex cover $C_{1}$ of $G$ such that $C\triangle C_{1}=\{t_{k},t_{\ell}\}$, where $C\triangle C_{1}:=(C\setminus C_{1})\cup(C_{1}\setminus C)$ is the symmetric difference.
Proof.
(a) $\Rightarrow$ (b) Letting $C_{1}=(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$, we show that $C_{1}$ is a vertex cover of $G$. Take $e=\{t_{i},t_{j}\}$. We may assume $t_{\ell}\notin e$, and $t_{i}\in C$ because $C$ is a vertex cover of $G$. Hence, if $t_{i}\neq t_{k}$, then $e\cap C_{1}\neq\emptyset$. Thus, we may assume $t_{i}=t_{k}$. Then, $t_{j}\in N_{G}(t_{k})$, and consequently $t_{j}\in C$ because $N_{G}(t_{k})\setminus C=\{t_{\ell}\}$. Hence, $t_{j}\in C\setminus\{t_{k}\}$ because $t_{i}\neq t_{j}$. Thus, $e\cap C_{1}\neq\emptyset$.
(b) $\Rightarrow$ (c) Letting $C_{1}=(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$, one has $C\setminus C_{1}=\{t_{k}\}$ and $C_{1}\setminus C=\{t_{\ell}\}$. Thus, the symmetric difference $C\triangle C_{1}$ is equal to $\{t_{k},t_{\ell}\}$.
(c) $\Rightarrow$ (a) If $t_{k}\in C\setminus C_{1}$ (resp. $t_{\ell}\in C\setminus C_{1}$), then $t_{\ell}\in C_{1}\setminus C$ (resp. $t_{k}\in C_{1}\setminus C$) because $C_{1}$ (resp. $C$) is a vertex cover of $G$. Thus, permuting $t_{k}$ and $t_{\ell}$ if necessary, we may assume that $t_{k}\in C\setminus C_{1}$. Clearly, $\{t_{\ell}\}\subset N_{G}(t_{k})\setminus C$. To show the other inclusion take $t_{i}\in N_{G}(t_{k})\setminus C$. Then, $\{t_{i},t_{k}\}\in E(G)$, and $t_{i}\in C_{1}$ because $t_{k}\notin C_{1}$ and $C_{1}$ is a vertex cover of $G$. Thus, $t_{i}\in C_{1}\setminus C$, and $t_{i}=t_{\ell}$ because $t_{i}\neq t_{k}$. ∎
Lemma 3.3.
Let $I_{c}(G)$ be the ideal of covers of a graph $G$ and let $\alpha_{e}(G)$ be its exchange number. The following hold.
  1. (a)
    (cf. [17, Theorem 3.5], [26, Remark 2.5]) ${\rm v}(I_{c}(G))=\alpha_{e}(G)-1$.
  2. (b)
    [17, Lemma 3.16] ${\rm v}(I_{c}(G))\geq\alpha_{0}(G)-1$.
  3. (c)
    ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$ if and only if there exists a vertex cover $C$ of $G$ with $\alpha_{0}(G)$ vertices satisfying the exchange property.
Proof.
(a) By Theorem 2.1, there exist an associated prime $(t_{k},t_{\ell})$ of $I_{c}(G)$, $\{t_{k},t_{\ell}\}\in E(G)$, and a set of vertices $A\subset V(G)$ such that
$(I_{c}(G)\colon t_{A})=(t_{k},\,t_{\ell})$
and ${\rm v}(I_{c}(G))=\deg(t_{A})$, where $t_{A}=\prod_{t_{i}\in A}t_{i}$. Then, $t_{k}t_{A}$ and $t_{\ell}t_{A}$ are in $I_{c}(G)$, and consequently $C=A\cup\{t_{k}\}$ and $C_{1}=A\cup\{t_{\ell}\}$ are vertex covers of $G$. Note that $\{t_{k},t_{\ell}\}\cap A=\emptyset$ because $I_{c}(G)$ is a squarefree monomial ideal and $t_{A}\notin I_{c}(G)$. Hence, $C\setminus C_{1}=\{t_{k}\}$ and $C_{1}\setminus C=\{t_{\ell}\}$ and, by Proposition 3.2, $C$ has the exchange property. Thus,
${\rm v}(I_{c}(G))=\deg(t_{A})=|C|-1\geq\alpha_{e}(G)-1,$
and consequently ${\rm v}(I_{c}(G))\geq\alpha_{e}(G)-1$. To show the reverse inequality note that, by definition of $\alpha_{e}(G)$, there is a vertex cover $C$ of $G$ satisfying the exchange property and $\alpha_{e}(G)=|C|$. Hence, there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$. Letting $A:=C\setminus\{t_{k}\}$, the sets $A\cup\{t_{k}\}$ and $A\cup\{t_{\ell}\}$ are minimal vertex covers of $G$. Thus, $t_{k}t_{A}\in I_{c}(G)$ and $t_{\ell}t_{A}\in I_{c}(G)$, that is, $\{t_{k},t_{\ell}\}\subset(I_{c}(G)\colon t_{A})$. From the embedding
$S/(I_{c}(G)\colon t_{A})\xhookrightarrow{\ \ t_{A}\ \ }S/I_{c}(G),$
and observing that any associated prime $\mathfrak{p}$ of $I_{c}(G)$ is of the form $\mathfrak{p}=(e)$ for some $e\in E(G)$ it follows that $(t_{k},t_{\ell})$ is an associated prime of $(I_{c}(G)\colon t_{A})$ and this is the only associated prime of $(I_{c}(G)\colon t_{A})$. Thus, $(I_{c}(G)\colon t_{A})=(t_{k},t_{\ell})$,
$\alpha_{e}(G)=|C|=\deg(t_{A})+1\geq{\rm v}(I_{c}(G))+1,$
and consequently $\alpha_{e}(G)-1\geq{\rm v}(I_{c}(G))$.
(b) This follows from part (a) noticing that $\alpha_{e}(G)\geq\alpha_{0}(G)$.
(c) The result follows from part (a) by noticing that $\alpha_{e}(C)=\alpha_{0}(G)$ if and only if there is a vertex cover $C$ of $G$ with $\alpha_{0}(G)$ vertices satisfying the exchange property. ∎
Let $J:=I_{c}(G)$ be the ideal of covers of an unmixed graph $G$, let $E(G^{\vee})=\{C_{i}\}_{i=1}^{r}$, and let $\mathcal{G}_{J}$ be the graph of $J$. From Section 2, recall that $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{J}$ if $|C_{i}\cap C_{j}|=|C_{i}|-1$. For all $C_{i},C_{j}$ in $V(\mathcal{G}_{J})$, $i\neq j$, let $\mathcal{G}_{J}^{(C_{i},C_{j})}$ denote the induced subgraph of $\mathcal{G}_{J}$ with vertex set
$V(\mathcal{G}_{J}^{(C_{i},C_{j})})=\{C_{k}\mid C_{k}\subset C_{i}\cup C_{j}\}.$
Proposition 3.4.
Let $G$ be an unmixed graph. The following conditions are equivalent.
  1. (a)
    The ideal $J$ is linearly presented.
  2. (b)
    $\mathcal{G}_{J}^{(C_{i},C_{j})}$ is connected for all $C_{i},C_{j}\in E(G^{\vee})$, $i\neq j$.
  3. (c)
    For all $C_{i},C_{j}\in E(G^{\vee})$, $i\neq j$, there is a path in $\mathcal{G}_{J}^{(C_{i},C_{j})}$ connecting $C_{i}$ and $C_{j}$ whose vertices are contained in $C_{i}\cup C_{j}$.
Proof.
It follows readily from Proposition 2.11. ∎
Theorem 3.5.
Let $G$ be an unmixed graph, let $C_{1},\ldots,C_{r}$ be the minimal vertex covers of $G$, and let $\mathcal{G}_{J}$ be the graph of $J=I_{c}(G)$. The following hold.
  1. (a)
    $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{J}$ if and only if there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C_{i}$, $t_{\ell}\notin C_{i}$ and $(C_{i}\setminus\{t_{k}\})\cup\{t_{\ell}\}=C_{j}$. In particular, non-isolated vertices of $\mathcal{G}_{J}$ have the exchange property.
  2. (b)
    $\mathcal{G}_{J}$ is not a discrete graph if and only if ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
  3. (c)
    If $I_{c}(G)$ is linearly presented, then all minimal vertex covers of $G$ have the exchange property and ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
Proof.
(a) $\Rightarrow$) Let $\{C_{i},C_{j}\}$ be an edge of $G$ and let $d=\alpha_{0}(G)$. Then, by definition of $\mathcal{G}_{J}$, one has $|C_{i}\cap C_{j}|=|C_{i}|-1=|C_{j}|-1=d-1$. Thus, we may assume
$C_{i}\cap C_{j}=\{t_{1},\ldots,t_{d-1}\},\ C_{i}=\{t_{1},\ldots,t_{d-1},t_{k}\},\ C_{j}=\{t_{1},\ldots,t_{d-1},t_{\ell}\},$
where $t_{k}\neq t_{\ell}$ and $t_{\ell}\notin C_{i}$. Then, $(C_{i}\setminus\{t_{k}\})\cup\{t_{\ell}\}=C_{j}$. We now show that $\{t_{k},t_{\ell}\}$ is an edge of $G$. By the minimality of the vertex cover $C_{i}$, there is $e\in E(G)$ such that $e\cap\{t_{1},\ldots,t_{d-1}\}=\emptyset$. As $C_{i}$ and $C_{j}$ are vertex covers of $G$, we get $e=\{t_{k},t_{\ell}\}$.
$\Leftarrow$) Let $C_{i},C_{j}$ be vertex covers of $G$ and let $\{t_{k},t_{\ell}\}$ be an edge of $G$ such that $t_{k}\in C_{i}$, $t_{\ell}\notin C_{i}$ and $(C_{i}\setminus\{t_{k}\})\cup\{t_{\ell}\}=C_{j}$. Then, using the inclusion-exclusion formula, from the equality
$(C_{i}\cup C_{j})\setminus\{t_{k},t_{\ell}\}=C_{i}\cap C_{j},$
we get $(|C_{i}|+|C_{j}|-|C_{i}\cap C_{j}|)-2=|C_{i}\cap C_{j}|$ and since $d=|C_{i}|=|C_{j}|$, one has
$|C_{i}\cap C_{j}|=d-1=|C_{i}|-1=|C_{j}|-1,$
that is, $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{J}$.
(b) $\Rightarrow$) Pick an edge $\{C_{i},C_{j}\}$ of $\mathcal{G}_{J}$. Then, by part (a), $C_{i}$ has the exchange property. Hence, by Lemma 3.3(c), we get ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
$\Leftarrow$) By Lemma 3.3(c), there is a vertex cover $C$ of $G$ with $\alpha_{0}(G)$ vertices satisfying the exchange property. Then, by part (a), $C$ is not an isolated vertex of $\mathcal{G}_{J}$.
(c) The graph $G$ has at least two minimal vertex covers because $G$ has at least one edge, that is, the graph $\mathcal{G}_{J}$ has at least two vertices. By Proposition 3.4, the graph $\mathcal{G}_{J}$ is connected. Then, by part (a), all minimal vertex covers of $G$ have the exchange property. Hence, by part (b), we obtain the equality ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$. ∎
Proposition 3.6.
Let $G$ be a graph. If $N_{G}(t_{k})\subset N_{G}[t_{\ell}]$ for some $\{t_{k},t_{\ell}\}\in E(G)$, then there exists a minimal vertex cover $C$ of $G$ such that $N_{G}(t_{k})\setminus C=\{t_{\ell}\}$.
Proof.
Pick a minimal vertex cover $C$ of $G$ that does not contain $t_{\ell}$. Then, $t_{k}\in C$. The inclusion $N_{G}(t_{k})\setminus C\supset\{t_{\ell}\}$ is clear. To show the reverse inclusion take $t_{i}\in N_{G}(t_{k})\setminus C$, then $t_{i}\in N_{G}(t_{k})$, and consequently $t_{i}\in N_{G}[t_{\ell}]$. Note that $N_{G}(t_{\ell})\subset C$ because $t_{\ell}\notin C$. Hence, if $t_{i}\neq t_{\ell}$, then $t_{i}\in C$, a contradiction. Thus, $t_{i}=t_{\ell}$, as required. ∎
Corollary 3.7.
Let $G$ be a graph and let $I_{c}(G)$ be its ideal of covers. The following hold.
  1. (a)
    Suppose $I(G)$ is sequentially Cohen–Macaulay. Then,
    ${\rm reg}(S/I_{c}(G))={\rm bight}(I(G))-1\geq{\rm v}(I_{c}(G))=\alpha_{e}(G)-1\geq\alpha_{0}(G)-1,$
    with equality everywhere if and only if $G$ is unmixed.
  2. (b)
    [29] $S/I(G)$ is Cohen–Macaulay if and only if ${\rm reg}(S/I_{c}(G))=\alpha_{0}(G)-1$.
  3. (c)
    [26, Corollary 3.9] If $S/I(G)$ is Cohen–Macaulay, then ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
Proof.
(a) The equality ${\rm reg}(S/I_{c}(G))={\rm bight}(I(G))-1$ was proved in [23, Theorem 3.31]. Then, by Eq. (1.1) and Lemma 3.3(a), we get
${\rm reg}(S/I_{c}(G))={\rm bight}(I(G))-1\geq{\rm v}(I_{c}(G))=\alpha_{e}(G)-1\geq\alpha_{0}(G)-1,$
and equality holds everywhere if and only if ${\rm bight}(I(G))=\alpha_{0}(G)$, and the latter equality holds if and only if $G$ is unmixed.
(b) By Theorem 2.4(b) and duality, one has ${\rm reg}(S/I_{c}(G))={\rm pd}_{S}(S/I(G))-1$. Hence, the result follows recalling that $\alpha_{0}(G)$ is the height of $I(G)$.
(c) The result follows from part (a) by recalling that $S/I(G)$ is Cohen–Macaulay if and only if $S/I(G)$ is sequentially Cohen–Macaulay and $I(G)$ is unmixed [33, Corollary 6.3.31]. ∎
For use below, let $\mathcal{F}$ be the set of all maximal stable set of $G^{\vee}$ and let $\mathcal{A}$ be the set of all $A$ such that $A$ is a stable set of $G^{\vee}$ and $N_{G^{\vee}}(A)$ is a minimal vertex cover of $G^{\vee}$.
Proposition 3.8.
Let $G$ be a graph. Then, ${\rm v}(I_{c}(G))=\dim(S/I_{c}(G))$ if and only if $\mathcal{F}=\mathcal{A}$.
Proof.
This follows by applying [17, Corollary 3.7] to the clutter $G^{\vee}$ and noticing that $G^{\vee}$ is unmixed because the minimal covers of $G^{\vee}$ are precisely the edges of $G$. ∎
Proposition 3.9.
Let $G$ be a graph with $s$ vertices. The following conditions are equivalent.
(a) $G$ is a codi-graph. (b) ${\rm depth}(S/I(G))=1$. (c) ${\rm reg}(S/I_{c}(G))=s-2$.
Proof.
(a)$\Leftrightarrow$(b) Let $\Delta$ be the independence complex of $G$ consisting of all stable sets of $G$ and let $K[\Delta]$ be its Stanley–Reisner ring. Note that $K[\Delta]=S/I(G)$. By [28], one has
[…]
where $\Delta^{i}=\{F\in\Delta\,|\,\dim(F)\leq i\}$ is the $i$-skeleton of $\Delta$ and $-1\leq i\leq\dim(\Delta)$. Note that $\Delta^{1}$ is the 1-skeleton of $\Delta^{i}$ for all $1\leq i\leq\dim(\Delta)$. As the facets of $\Delta^{1}$ are the edges of $\overline{G}$ and $\Delta^{1}$ has dimension $1$, by [33, Corollary 6.3.14], $K[\Delta^{1}]$ is Cohen–Macaulay if and only if $\overline{G}$ is connected. Using the fact that the 1-skeleton of a Cohen–Macaulay complex is Cohen–Macaulay [33, Proposition 6.3.17], we obtain that $K[\Delta^{i}]$ is not Cohen–Macaulay for all $1\leq i\leq\dim(\Delta)$ if and only $\Delta^{1}$ is not Cohen–Macaulay. Hence, by Eq. (3.1), ${\rm depth}(S/I(G))=1$ if and only if $\overline{G}$ is disconnected.
(b)$\Leftrightarrow$(c) From Auslander–Buchsbaum and Terai’s formulas (Eq. (2.1), Theorem 2.4(b)):
${\rm reg}(S/I_{c}(G))={\rm pd}_{S}(S/I(G))-1,\quad{\rm pd}_{S}(S/I(G))+{\rm depth}(S/I(G))=s,$
we obtain ${\rm reg}(S/I_{c}(G))=s-{\rm depth}(S/I(G))-1$, and the equivalence of (b) and (c) follows. ∎
Theorem 3.10.
Let $G$ be a graph with $s$ vertices such that $E(G)\neq\emptyset$ and let $I_{c}(G)$ be its ideal of covers. The following conditions are equivalent.
  1. (a)
    ${\rm v}(I_{c}(G))=s-2$.
  2. (b)
    ${\rm v}(I_{c}(G))={\rm reg}(S/I_{c}(G))=\dim(S/I_{c}(G))=s-2$ and $G$ is a codi-graph.
  3. (c)
    Every vertex cover of $G$ with the exchange property has exactly $s-1$ vertices.
  4. (d)
    ${\rm v}(I_{c}(H))=|V(H)|-2$ for any induced subgraph $H$ of $G$ with $E(H)\neq\emptyset$.
  5. (e)
    Any induced subgraph $H$ of $G$ with $E(H)\neq\emptyset$ is connected.
  6. (f)
    $G$ is a complete multipartite graph.
Proof.
(a)$\Rightarrow$(b) As ${\rm v}(I_{c}(G))=s-2$, equality holds everywhere in Eq. (1.1). That $G$ is a codi-graph follows from Proposition 3.9.
(b)$\Rightarrow$(c) Take any vertex cover $C$ of $G$ with the exchange property. Then, there exists $e=\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$. Letting $A=C\setminus\{t_{k}\}$, note that $A$ is a stable set of $G^{\vee}$ because $e\in E(G)$ and $e\cap A=\emptyset$. As $A\cup\{t_{k}\}$ and $A\cup\{t_{\ell}\}$ are vertex covers of $G$, one has $N_{G^{\vee}}(A)\supset e$, and consequently $(I_{c}(G)\colon t_{A})\supset(e)$. Hence, from the proof of the second part of Lemma 3.3(a), we obtain $(I_{c}(G)\colon t_{A})=(e)$. Therefore, $N_{G^{\vee}}(A)=e$ and consequently $A\in\mathcal{A}$. Then, by Proposition 3.8, $\mathcal{F}=\mathcal{A}$ and $A\in\mathcal{F}$. Thus, $A$ is a maximal stable set of $G^{\vee}$ and $V(G)\setminus A$ is a minimal vertex cover of $G^{\vee}$. Since the minimal vertex covers of $G^{\vee}$ are the edges of the graph $G$, we get
$12=|V(G)\setminus A|=|V(G)|-|A|=s-(|C|-1)\ \therefore\ |C|=s-$ 1 .
(c)$\Rightarrow$(a) By Lemma 3.3(a), there exists a vertex cover $C$ of $G$ with the exchange property such that ${\rm v}(I_{c}(G))=|C|-1$. Thus, ${\rm v}(I_{c}(G))=s-2$.
Thus, we have proved that (a), (b), and (c) are equivalent conditions.
(a)$\Rightarrow$(d) Let $H$ be an induced subgraph of $G$ with $E(H)\neq\emptyset$ and let $D$ be any vertex cover of $H$ with the exchange property. Then, there is $\{t_{k},t_{\ell}\}\in E(H)$, $t_{k}\in D$, $t_{\ell}\notin D$, such that $(D\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $H$. Letting $C=D\cup(V(G)\setminus V(H))$, note that $C$ is a vertex cover of $G$. From the equality
$(C\setminus\{t_{k}\})\cup\{t_{\ell}\}=((D\setminus\{t_{k}\})\cup\{t_{\ell}\})\cup(V(G)\setminus V(H)),$
we obtain that $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is also a vertex cover of $G$. Thus, $C$ has the exchange property relative to $G$. Then, applying the equivalence of (a) and (c), we get $s-2=|C|-1=|D|+(s-n)-1$, where $n=|V(H)|$, and consequently $|D|=n-1$. Then, applying the equivalence of (a) and (c) to the graph $H$, we get that ${\rm v}(I_{c}(H))=n-2$.
(d)$\Rightarrow$(e) Applying the equivalence of (a) and (b) to the graph $H$, we obtain that the complement $\overline{H}$ of $H$ is a disconnected graph. Then, $H$ is a connected graph.
(e)$\Rightarrow$(c) Let $C$ be any vertex cover of $G$ with the exchange property. Then, by Proposition 3.2, there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $N_{G}(t_{k})\setminus C=\{t_{\ell}\}$. It suffices to show that $C\cup\{t_{\ell}\}=V(G)$. We argue by contradiction assuming that there is $t_{i}\in V(G)$ and $t_{i}\notin C\cup\{t_{\ell}\}$. Since the induced subgraph $G[\{t_{k},t_{\ell},t_{i}\}]$ is connected, either $\{t_{i},t_{k}\}\in E(G)$ or $\{t_{i},t_{\ell}\}\in E(G)$. If $\{t_{i},t_{k}\}\in E(G)$, then $t_{i}\in N_{G}(t_{k})\setminus C=\{t_{\ell}\}$, a contradiction. If $\{t_{i},t_{\ell}\}\in E(G)$, then $t_{i}\in C$ because $C$ is a vertex cover of $G$, a contradiction.
Thus, we have proved that conditions (a) to (e) are equivalent. Hence, to prove that (a) and (f) are equivalent, we need only show that (a) implies (f) and (f) implies (e).
(a)$\Rightarrow$(f) Let $H_{1},\ldots,H_{r}$ be the connected components of $\overline{G}$ and let $G_{i}=\overline{H_{i}}$ be its complement for $i=1,\ldots,r$. Then, $G_{1},\ldots,G_{r}$ are pairwise disjoint graphs and $G=G_{1}*\cdots*G_{r}$. If $r=1$, by the equivalence (a) and (b), $\overline{G}=H_{1}$ is disconnected, a contradiction. Thus, $r\geq 2$. To show that $G$ is multipartite we show that $G_{i}$ is a discrete graph for all $i$. We argue by contradiction assuming that $E(G_{i})\neq\emptyset$ for some $i$. As $G_{i}$ is an induced subgraph of $G$, by applying the equivalence of (a), (d), and (b) to the graph $G_{i}$, we obtain that $G_{i}$ is a connected codi-graph, and consequently $\overline{G_{i}}=H_{i}$ is disconnected, a contradiction.
(f)$\Rightarrow$(e) Let $G=G_{1}*\cdots*G_{r}$ be a complete multipartite graph and let $H$ be an induced subgraph of $G$ with $E(H)\neq\emptyset$. Pick an edge $\{t_{i},t_{j}\}\in E(H)$. For any $t_{k}\in V(H)\setminus\{t_{i},t_{j}\}$, one has that either $t_{k}\in N_{H}(t_{i})$ or $t_{k}\in N_{H}(t_{j})$ because $H$ is an induced subgraph of $G$. Thus, $H$ is a connected graph. ∎
Corollary 3.11.
If $G$ is a graph with $s$ vertices and ${\rm v}(I_{c}(G))=s-2$, then
  1. (a)
    $I_{c}(G)$ is Cohen–Macaulay,
  2. (b)
    ${\rm v}_{\mathfrak{p}}(I_{c}(G))=s-2$ for all $\mathfrak{p}\in{\rm Ass}(I_{c}(G))$, and
  3. (c)
    $(I_{c}(G)\colon\mathfrak{p})/I_{c}(G)$ is a principal ideal of $S/I_{c}(G)$ for all $\mathfrak{p}\in{\rm Ass}(I_{c}(G))$.
Proof.
(a) By Theorem 3.10, $G$ is a multipartite graph. Hence, $\overline{G}$ is a disjoint union of complete graphs. Thus, $\overline{G}$ is chordal and, by a result of Lyubeznik [19], $I_{c}(G)$ is Cohen–Macaulay.
(b) Let $\mathfrak{p}\in{\rm Ass}(I_{c}(G))$. Then, $\mathfrak{p}=(t_{k},t_{\ell})$ for some $\{t_{k},t_{\ell}\}\in E(G)$ since $I_{c}(G)=\bigcap_{e\in E(G)}(e)$. Pick $A\subset V(G)$ such that $(I_{c}(G)\colon t_{A})=\mathfrak{p}$ and $\deg(t_{A})={\rm v}_{\mathfrak{p}}(I_{c}(G))$. Note that $A\cap\{t_{k},t_{\ell}\}=\emptyset$ because $I_{c}(G)$ is squarefree. Setting $C=A\cup\{t_{k}\}$, it follows that $C$ is a vertex cover of $G$ with the exchange property. Hence, by Theorem 3.10, $|C|=s-1$. Thus, ${\rm v}_{\mathfrak{p}}(I_{c}(G))=|A|=|C|-1=s-2$.
(c) Let $t_{A}+I_{c}(G)$, $A\subset V(G)$, be a minimal generator of $N:=(I_{c}(G)\colon\mathfrak{p})/I_{c}(G)$ and let $\{t_{k},t_{\ell}\}$ be an edge of $G$ such that $\mathfrak{p}=(t_{k},t_{\ell})$. Note that $A\cap\{t_{k},t_{\ell}\}=\emptyset$ because $I_{c}(G)$ is squarefree. Then, $A\cup\{t_{k}\}$ and $A\cup\{t_{\ell}\}$ are vertex covers of $G$ with the exchange property. Hence, by Theorem 3.10, one has $A\cup\{t_{k},t_{\ell}\}=V(G)$, and consequently any other minimal generator of $N$ is equal to $t_{A}+I_{c}(G)$. ∎
Proposition 3.12.
Let $G=G_{1}*G_{2}$ be the join of two vertex disjoint graphs $G_{1}$ and $G_{2}$. If $G$ is not a complete multipartite graph, then
  1. (a)
    ${\rm v}(I_{c}(G))=\min\{{\rm v}(I_{c}(G_{1}))+|V(G_{2})|,\,{\rm v}(I_{c}(G_{2}))+|V(G_{1})|\}$, and
  2. (b)
    ${\rm v}(I_{c}(G))={\rm v}(I_{c}(G_{1}))+|V(G_{2})|$ if $G_{2}$ is a discrete graph.
Proof.
(a) We set $V_{i}=V(G_{i})$, $i=1,2$, and $s=|V(G)|$. First we show the inequalities:
(3.2) $\alpha_{e}(G)\leq\begin{cases}\alpha_{e}(G_{1})+|V_{2}|&\mbox{if }E(G_{1})\neq\emptyset,\\ \alpha_{e}(G_{2})+|V_{1}|&\mbox{if }E(G_{2})\neq\emptyset.\end{cases}$
Assume that $E(G_{1})\neq\emptyset$. By Lemma 3.3, there is a vertex cover $D$ of $G_{1}$ with the exchange property such that $|D|=\alpha_{e}(G_{1})$. Then, there exists $\{t_{i},t_{j}\}\in E(G_{1})$ such that $t_{i}\in D$, $t_{j}\notin D$, and $(D\setminus\{t_{i}\})\cup\{t_{j}\}$ is a vertex cover of $G_{1}$. Then,
$D\cup V_{2}\ \mbox{ and }\ ((D\cup V_{2})\setminus\{t_{i}\})\cup\{t_{j}\}$
are vertex covers of $G$. Thus, $D\cup V_{2}$ is a vertex cover of $G$ with the exchange property, and
$\alpha_{e}(G_{1})+|V_{2}|=|D|+|V_{2}|=|D\cup V_{2}|\geq\alpha_{e}(G).$
If $E(G_{2})\neq\emptyset$, by a similar argument, we obtain $\alpha_{e}(G)\leq\alpha_{e}(G_{2})+|V_{1}|$.
There is a vertex cover $C$ of $G$ with the exchange property such that $|C|=\alpha_{e}(G)$ (Lemma 3.3). Then, there exists $\{t_{k},t_{\ell}\}\in E(G)$ such that $t_{k}\in C$, $t_{\ell}\notin C$, and $(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ is a vertex cover of $G$. We claim that $\{t_{k},t_{\ell}\}\in E(G_{i})$ for some $i\in\{1,2\}$. We argue by contradiction assuming that $t_{k}\in V_{1}$ and $t_{\ell}\in V_{2}$ (the case $t_{\ell}\in V_{1}$ and $t_{k}\in V_{2}$ can be treated similarly). If $|C|=s-1$, then ${\rm v}(I_{c}(G))=s-2$ and, by Theorem 3.10, $G$ is a complete multipartite graph, a contradiction. Thus, $|C|\leq s-2$ and $C\subsetneq V(G)\setminus\{t_{\ell}\}$. Pick $t_{i}\in V(G)\setminus\{t_{\ell}\}$, $t_{i}\notin C$. Noticing that $V_{1}\subset C$ because $t_{\ell}\in V_{2}\setminus C$ and $G=G_{1}*G_{2}$, we get that $t_{i}\notin V_{1}$ and $t_{i}\in V_{2}$. Then, $\{t_{i},t_{k}\}\in E(G)$ and consequently $t_{i}\in(C\setminus\{t_{k}\})\cup\{t_{\ell}\}$ because the latter is a vertex cover of $G$, a contradiction. This proves that $\{t_{k},t_{\ell}\}\in E(G_{i})$ for some $i\in\{1,2\}$.
We may assume $\{t_{k},t_{\ell}\}\in E(G_{1})$ (the case $\{t_{k},t_{\ell}\}\in E(G_{2})$ can be treated similarly). Note that $C\supset V_{2}$ because $t_{\ell}\notin C$ and $t_{\ell}\in V_{1}$. Then,
$C\setminus V_{2}\ \mbox{ and }\ ((C\setminus V_{2})\setminus\{t_{k}\})\cup\{t_{\ell}\}$
are vertex covers of $G_{1}$, $C\setminus V_{2}$ is a vertex cover of $G_{1}$ with the exchange property, and
(3.3) $\alpha_{e}(G_{1})\leq|C\setminus V_{2}|=|C|-|V_{2}|=\alpha_{e}(G)-|V_{2}|.$
Hence, from Eqs. (3.2)-(3.3), we get $\alpha_{e}(G)=\alpha_{e}(G_{1})+|V_{2}|$. Furthermore, if $E(G_{2})\neq\emptyset$, by Eq. (3.2), one has $\alpha_{e}(G)\leq\alpha_{e}(G_{2})+|V_{1}|$ and
$\alpha_{e}(G)=\min\{\alpha_{e}(G_{1})+|V_{2}|,\,\alpha_{e}(G_{2})+|V_{1}|\}.$
Hence, by Lemma 3.3, we obtain
${\rm v}(I_{c}(G))=\min\{{\rm v}(I_{c}(G_{1}))+|V_{2}|,\,{\rm v}(I_{c}(G_{2}))+|V_{1}|\}.$
(b) If $E(G_{2})=\emptyset$, by the proof of part (a) it follows that ${\rm v}(I_{c}(G))={\rm v}(I_{c}(G_{1}))+|V_{2}|$. ∎
Corollary 3.13.
Let $G_{1}$ be the vertex disjoint union of $k+1$ complete graphs with two vertices, let $G_{2}$ be a discrete graph with one vertex, and let $G=G_{1}*G_{2}$ be their join. Then,
${\rm reg}(S/I_{c}(G))-{\rm v}(I_{c}(G))=k.$
Proof.
The edge ideal $I(G_{1})$ is a complete intersection, hence it is Cohen–Macaulay. Then, by Corollary 3.7 and Proposition 3.12, we obtain
${\rm v}(I_{c}(G_{1}))=\alpha_{0}(G_{1})-1=k\ \mbox{ and }\ {\rm v}(I_{c}(G))={\rm v}(I_{c}(G_{1}))+1=k+1.$
As $G$ is a codi-graph, by Proposition 3.9, one has ${\rm reg}(S/I_{c}(G))=|V(G)|-2=2k+1$. Hence, the difference of the regularity of $S/I_{c}(G)$ and the v-number of $I_{c}(G)$ is equal to $k$. ∎
Proposition 3.14.
Let $G$ be a complete $k$-partite graph. The following hold.
  1. (a)
    If $k=2$, then each edge of $G$ satisfies the (P) property.
  2. (b)
    If $k\geq 3$, then no edge of $G$ satisfies the (P) property.
  3. (c)
    $G$ is unmixed if and only if $G$ is a homogeneous multipartite graph.
Proof.
We can write $G=G_{1}*\cdots*G_{k}$, where $G_{i}$ is a discrete graph for $1\leq i\leq k$.
(a) Take $e\in E(G)$, then $e=\{t_{1},t_{2}\}$ with $t_{1}\in V(G_{1})$ and $t_{2}\in V(G_{2})$ since $k=2$. Now, let $\{t_{1},t^{\prime}_{1}\},\{t_{2},t^{\prime}_{2}\}$ be two edges of $G$ distinct from $e$, then $t^{\prime}_{1}\in V(G_{2})$ and $t^{\prime}_{2}\in V(G_{1})$. Hence, $\{t^{\prime}_{1},t^{\prime}_{2}\}\in E(G)$, since $G$ is a complete $2$-partite graph. Thus, $e$ satisfies the (P) property.
(b) Take $e\in E(G)$, then $e=\{t_{1},t_{2}\}$ with $t_{1}\in V(G_{i_{1}})$ and $t_{2}\in V(G_{i_{2}})$ where $1\leq i_{1}<i_{2}\leq k$, since $G$ is $k$-partite. Since $k\geq 3$, there is $t_{3}\in V(G_{i_{3}})$ with $i_{3}\notin\{i_{1},i_{2}\}$. Thus, $\{t_{1},t_{3}\}\in E(G)$ and $\{t_{2},t_{3}\}\in E(G)$, since $G$ is a complete $k$-partite graph. But $\{t_{3},t_{3}\}\notin E(G)$, since $G$ is a simple graph. Hence, $e$ does not satisfy the (P) property.
(c) Note that the maximal stable sets of $G$ are $V(G_{1}),\ldots,V(G_{k})$. Hence, $G$ is unmixed if and only if $|V(G_{1})|=\cdots=|V(G_{k})|$, i.e., if and only if $G$ is homogeneous. ∎
Corollary 3.15.
Let $G$ be a homogeneous complete $k$-partite graph. The following conditions are equivalent.
${\rm(a)}\ \mathcal{G}_{I_{c}(G)}\mbox{ is a complete graph}.\quad{\rm(b)}\ G\mbox{ is a complete graph}.\quad{\rm(c)}\ {\rm v}(I_{c}(G))=\alpha_{0}(G)-1.$
Proof.
Let $s=|V(G)|$. We can write $G=G_{1}*\cdots*G_{k}$, where $G_{i}$ is a discrete graph with $p$ vertices for $i=1,\ldots,k$. Then, $G$ is unmixed by Proposition 3.14 and ${\rm v}(I_{c}(G))=s-2$ by Theorem 3.10. Then, $\alpha_{0}(G)=s-p$ because all maximal stable sets of $G$ have $p$ elements.
(b)$\Leftrightarrow$(c) The equality ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$ holds if and only if $s-2=s-p-1$, i.e., if and only if $p=1$, that is, if and only if $G$ is a complete graph.
(a)$\Rightarrow$(c) As $\mathcal{G}_{I_{c}(G)}$ is connected, by Theorem 3.5(b), ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
(b)$\Rightarrow$(a) As $G$ is complete, $G$ is unmixed and $\alpha_{0}(G)=s-1$. Letting $V(G)=\{t_{1},\ldots,t_{s}\}$ and $C_{i}=V(G)\setminus\{t_{i}\}$ for $i=1,\ldots,s$, one has $E(G^{\vee})=\{C_{i}\}_{i=1}^{s}$. Since $|C_{i}\cap C_{j}|=|C_{i}|-1$ for $i\neq j$, $\{C_{i},C_{j}\}$ is an edge of $\mathcal{G}_{I_{c}(G)}$ for $i\neq j$, and $\mathcal{G}_{I_{c}(G)}$ is a complete graph. ∎
4. The graphs of ideals of covers
Let $e_{i},1\leq i\leq s$, be the $i$-th unit vector of ${\mathbb{R}}^{s}$, let $C_{1},\ldots,C_{r}$ be the minimal vertex covers of an unmixed graph $G$ with $d=\alpha_{0}(G)$, let $\{{t_{c_{1}}}\}_{i=1}^{r}$ be the minimal set of generators of the ideal of covers $I_{c}(G)$ of $G$, where $t_{c_{i}}=\prod_{t_{j}\in C_{i}}t_{j}$, let $\mathcal{G}=\mathcal{G}_{I_{c}(G)}$ be the graph of $I_{c}(G)$, and let $\mathcal{K}_{r}$ be the complete graph with vertex set $V(\mathcal{G})$.
Proposition 4.1.
Let $G$ be an unmixed graph and let $\mathcal{LS}(I_{c}(G))$ be the linear syzygy matrix of $I_{c}(G)$. Then, $\mathcal{G}_{I_{c}(G)}$ is connected if and only if ${\rm rank}(\mathcal{LS}(I_{c}(G)))=r-1$.
Proof.
Let $c$ be the number of connected components of $\mathcal{G}_{I_{c}(G)}$. Then, by Proposition 2.14,
${\rm rank}(\mathcal{LS}(I_{c}(G)))=r-c,$
and consequently $\mathcal{G}_{I_{c}(G)}$ is connected if and only if ${\rm rank}(\mathcal{LS}(I_{c}(G)))=r-1$. ∎
Proposition 4.2.
Let $G$ be an unmixed graph and let $\{C_{1},C_{2},C_{3}\}$ be a path of length two of the graph $\mathcal{G}$ of $I_{c}(G)$. The following hold.
  1. (a)
    There exist $\epsilon_{i}=\{t_{k_{i}},t_{\ell_{i}}\}\in E(G)$, $i=1,2$, such that
    1. (1)
      $C_{2}=(C_{1}\setminus\{t_{k_{1}}\})\cup\{t_{\ell_{1}}\}=(C_{1}\cap C_{2})\cup\{t_{\ell_{1}}\}$, $t_{k_{1}}\in C_{1}$, $t_{\ell_{1}}\notin C_{1}$,
    2. (2)
      $C_{3}=(C_{2}\setminus\{t_{k_{2}}\})\cup\{t_{\ell_{2}}\}=(C_{2}\cap C_{3})\cup\{t_{\ell_{2}}\}$, $t_{k_{2}}\in C_{2}$, $t_{\ell_{2}}\notin C_{2}$,
    3. (3)
      $t_{\ell_{1}}\neq t_{k_{2}}$, and $C_{2}\subset C_{1}\cup C_{3}$.
  2. (b)
    If the path $\{C_{1},C_{2},C_{3}\}$ of $\mathcal{G}$ is induced, then $\epsilon_{1}\cap\epsilon_{2}=\emptyset$, $|C_{1}\cap C_{2}\cap C_{3}|=\alpha_{0}(G)-2$, $|C_{1}\cap C_{3}|=\alpha_{0}(G)-2$, and $t_{\ell_{2}}\notin C_{1}$.
Proof.
(a) The existence of $\epsilon_{1}$ and $\epsilon_{2}$ satisfying (1) and (2) follows from Theorem 3.5 and its proof. Now we prove part (3). To prove that $t_{\ell_{1}}\neq t_{k_{2}}$, we argue by contradiction assuming $t_{\ell_{1}}=t_{k_{2}}$. We claim that $C_{1}=C_{3}$. Since $|C_{1}|=|C_{3}|$, it suffices to show the inclusion $C_{1}\supset C_{3}$. Take $t_{i}\in C_{3}$. Note that $t_{\ell_{2}}\in C_{1}$ because $C_{1}$ is a vertex cover, $\{t_{k_{2}},t_{\ell_{2}}\}=\{t_{\ell_{1}},t_{\ell_{2}}\}\in E(G)$ and $t_{\ell_{1}}\notin C_{1}$. Thus, if $t_{i}=t_{\ell_{2}}$, one has $t_{i}\in C_{1}$. If $t_{i}\neq t_{\ell_{2}}$, by (2), we get $t_{i}\in C_{2}\setminus\{t_{k_{2}}\}$ and $t_{i}\neq t_{k_{2}}$. Then, by (1), $t_{i}\in C_{1}\setminus\{t_{k_{1}}\}$ or $t_{i}=t_{\ell_{1}}$. In the first case $t_{i}\in C_{1}$. In the second case $t_{i}=t_{\ell_{1}}=t_{k_{2}}$, a contradiction since $t_{i}\neq t_{k_{2}}$. Thus, $t_{i}\in C_{1}$. This proves the equality $C_{1}=C_{3}$, a contradiction because $C_{1},C_{2},C_{3}$ are distinct. To show the inclusion $C_{2}\subset C_{1}\cup C_{3}$ take $t_{i}\in C_{2}$. By (1), either $t_{i}\in C_{1}\setminus\{t_{k_{1}}\}\subset C_{1}$ or $t_{i}=t_{\ell_{1}}$. Thus, we may assume $t_{i}=t_{\ell_{1}}$. Recalling that $t_{\ell_{1}}\neq t_{k_{2}}$, we get $t_{i}=t_{\ell_{1}}\in C_{2}\setminus\{t_{k_{2}}\}\subset C_{3}$.
(b) To show that $\epsilon_{1}\cap\epsilon_{2}=\emptyset$ it suffices to show that the following conditions hold.
${\rm(i)}\ t_{\ell_{1}}\neq t_{\ell_{2}},\quad{\rm(ii)}\ t_{k_{1}}\neq t_{k_{2}},\quad{\rm(iii)}\ t_{\ell_{1}}\neq t_{k_{2}},\quad{\rm(iv)}\ t_{k_{1}}\neq t_{\ell_{2}}.$
(i) Assume that $t_{\ell_{1}}=t_{\ell_{2}}$. Then, by (1), $t_{\ell_{2}}=t_{\ell_{1}}\in C_{2}$, a contradiction because by the last part of (2) we have $t_{\ell_{2}}\notin C_{2}$.
(ii) Assume that $t_{k_{1}}=t_{k_{2}}$. Then, by (1)-(2), $t_{k_{1}}=t_{k_{2}}\in C_{1}\cap C_{2}=C_{1}\setminus\{t_{k_{1}}\}$, a contradiction.
(iii) This follows at once from part (a) item (3).
(iv) Assume that $t_{k_{1}}=t_{\ell_{2}}$. Note that $t_{k_{2}}\in C_{1}\cap C_{2}$. Indeed, by (2), $t_{k_{2}}\in C_{2}$ and, by (3), $t_{k_{2}}\neq t_{\ell_{1}}$. Then, by (1), $t_{k_{2}}\in C_{1}\setminus\{t_{k_{1}}\}\subset C_{1}$. From (2), we obtain
$C_{1}\cap C_{3}$ $=C_{1}\cap((C_{2}\setminus\{t_{k_{2}}\})\cup\{t_{\ell_{2}}\})=(C_{1}\cap(C_{2}\setminus\{t_{k_{2}}\}))\cup(C_{1}\cap\{t_{\ell_{2}}\})$
$=((C_{1}\cap C_{2})\setminus\{t_{k_{2}}\})\cup(C_{1}\cap\{t_{k_{1}}\})\quad\therefore$
$|C_{1}\cap C_{3}|$ $=|(C_{1}\cap C_{2})\setminus\{t_{k_{2}}\}|+|C_{1}\cap\{t_{k_{1}}\}|-|((C_{1}\cap C_{2})\setminus\{t_{k_{2}}\})\cap(C_{1}\cap\{t_{\ell_{2}}\})|$
$=((\alpha_{0}(G)-1)-1)+1-0=\alpha_{0}(G)-1=|C_{1}|-1=|C_{3}|-1.$
Hence, $\{C_{1},C_{3}\}$ is an edge of the graph $\mathcal{G}$ of $I_{c}(G)$, a contradiction because $\{C_{1},C_{2},C_{3}\}$ is an induced path of $\mathcal{G}$. Letting $d=\alpha_{0}(G)$, we now show that $|\bigcap_{i=1}^{3}C_{i}|=d-2$. By (3), $C_{2}\subset C_{1}\cup C_{3}$. Then, $C_{1}\cup C_{3}=C_{1}\cup C_{2}\cup C_{3}$ and, by the inclusion-exclusion principle (Lemma 2.15), one has
$|C_{1}\cup C_{3}|$ $=2d-|C_{1}\cap C_{3}|=|C_{1}\cup C_{2}\cup C_{3}|=3d-2(d-1)-|C_{1}\cap C_{3}|+|C_{1}\cap C_{2}\cap C_{3}|,$
and consequently, cancelling out $-|C_{1}\cap C_{3}|$, we obtain $|C_{1}\cap C_{2}\cap C_{3}|=d-2$. Finally, we show that $|C_{1}\cap C_{3}|$ is equal to $\alpha_{0}(G)-2$ and $t_{\ell_{2}}\notin C_{1}$. By (2) we obtain
$C_{1}\cap C_{3}$ $=C_{1}\cap((C_{2}\cap C_{3})\cup\{t_{\ell_{2}}\})=(C_{1}\cap C_{2}\cap C_{3})\cup(C_{1}\cap\{t_{\ell_{2}}\})\quad\therefore$
$|C_{1}\cap C_{3}|$ $=(d-2)+|C_{1}\cap\{t_{\ell_{2}}\}|-|C_{1}\cap C_{2}\cap C_{3}\cap\{t_{\ell_{2}}\}|=(d-2)+|C_{1}\cap\{t_{\ell_{2}}\}|.$
Then, $t_{\ell_{2}}\notin C_{1}$ because $\{C_{1},C_{3}\}$ is not an edge of $\mathcal{G}$. Thus, $|C_{1}\cap C_{3}|=d-2$. ∎
Proposition 4.3.
Let $G$ be an unmixed graph. A path $\{C_{1},\ldots,C_{n}\}$ of the graph $\mathcal{G}$ of $I_{c}(G)$ is a path of $\mathcal{G}^{(C_{1},C_{n})}$ if and only if
[…]
Proof.
$\Rightarrow$) If $\{C_{1},\ldots,C_{n}\}$ is a path of $\mathcal{G}^{(C_{1},C_{n})}$, then $C_{k}\subset C_{1}\cup C_{n}$ for all $1<k<n$. Hence,
$C_{k}=(C_{k}\cap C_{1})\cup(C_{k}\cap C_{n})\ \therefore\ \alpha_{0}(G)=|C_{k}\cap C_{1}|+|C_{k}\cap C_{n}|-|C_{1}\cap C_{k}\cap C_{n}|.$
$\Leftarrow$) By assumption $C_{k}$ and $(C_{k}\cap C_{1})\cup(C_{k}\cap C_{n})$ have the same cardinality for all $1<k<n$. Since the latter is contained $C_{k}$, one has $C_{k}=(C_{k}\cap C_{1})\cup(C_{k}\cap C_{n})$. Thus, $C_{k}\subset C_{1}\cup C_{n}$. ∎
Lemma 4.4.
If $t$ is not an isolated vertex of a graph $G$, then there are minimal vertex covers $C$ and $C^{\prime}$ of $G$ such that $t\in C^{\prime}$ and $t\notin C$.
Proof.
Since $t$ is not an isolated vertex of $G$, there is $t^{\prime}\in V(G)$ such that $\{t,t^{\prime}\}\in E(G)$. Then, there are maximal stable sets $A$ and $A^{\prime}$ such that $t\in A$ and $t^{\prime}\in A^{\prime}$. So, $t\notin A^{\prime}$, since $A^{\prime}$ is a stable set. Then, $C=V(G)\setminus A$ and $C^{\prime}=V(G)\setminus A^{\prime}$ are minimal vertex covers of $G$ such that $t\notin C$, and $t\in C^{\prime}$. ∎
Proposition 4.5.
Let $G$ be an unmixed graph without isolated vertices. If the graph $\mathcal{G}$ of $I_{c}(G)$ is connected, then $V(G)$ is the union of the exchange edges of $G$.
Proof.
Let $t\in V(G)$. By Lemma 4.4, there are minimal vertex covers $C$ and $C^{\prime}$ of $G$ such that $t\in C^{\prime}\setminus C$. As $\mathcal{G}$ is connected, there is a path $\mathcal{P}=\{C_{1}=C,C_{2},\ldots,C_{n}=C^{\prime}\}$ in $\mathcal{G}$. Let
$j=\mbox{min}\{i\mbox{ }|\mbox{ }C_{i}\in V(\mathcal{P})\mbox{ and }t\in C_{i}\},$
then $2\leq j\leq n$ since $t\notin C_{1}$ and $t\in C_{n}$. So, $t\in C_{j}\setminus C_{j-1}$. Furthermore, $\{C_{j-1},C_{j}\}$ is an edge of $\mathcal{G}$ and, by Theorem 3.5, $C_{j}=(C_{j-1}\setminus\{t^{\prime}\})\cup\{t\}$ for some $\{t,t^{\prime}\}\in E(G)$. Hence, $\{t,t^{\prime}\}$ is an exchange edge of $G$. Therefore, $V(G)$ is the union of the exchange edges. ∎
Proposition 4.6.
Let $G$ be an unmixed graph, let $\mathcal{G}$ be the graph of $I_{c}(G)$, and let $\{C_{1},\ldots,C_{n}\}$ be a path of $\mathcal{G}$. The following conditions are equivalent.
  1. (a)
    $C_{i}\subset C_{p}\cup C_{q}$ for all integers $1\leq p\leq i\leq q\leq n$.
  2. (b)
    $|\bigcap_{j=p}^{q}C_{j}|=\alpha_{0}(G)-(q-p)$ for all integers $1\leq p\leq q\leq n$.
Proof.
(a)$\Rightarrow$(b) We proceed by induction on $q$. The equality in (b) is clear if $q=p$ or $q=p+1$. Assume that $q>p+1$ and the equality in (b) valid for $q-1$. Consider the list
$C_{p},\,C_{p+1},\ldots,C_{q-2},\,C_{q-1},\,C_{q}.$
By hypothesis $C_{q-1}\subset C_{p}\cup C_{q},\,C_{q-1}\subset C_{p+1}\cup C_{q},\ldots,C_{q-1}\subset C_{q-2}\cup C_{q}$. Hence,
$C_{q-1}=(C_{q-1}\cap C_{q})\cup(C_{p}\cap C_{p+1}\cap\cdots\cap C_{q-2}\cap C_{q-1}).$
Then, setting $d=\alpha_{0}(G)$, by induction and applying inclusion-exclusion, one has
$d=(d-1)+\Bigg{|}\bigcap_{j=p}^{q-1}C_{j}\Big{|}-\Bigg{|}\bigcap_{j=p}^{q}C_{j}\Bigg{|}=(d-1)+(d-(q-1-p))-\Bigg{|}\bigcap_{j=p}^{q}C_{j}\Bigg{|}.$
Thus, $|\bigcap_{j=p}^{q}C_{j}|=d-(q-p)$ and the proof is complete.
(b)$\Rightarrow$(a) We proceed by induction on $i$. The inclusion $C_{i}\subset C_{p}\cup C_{q}$ is clear if $i=p$ or $i=q$. Assume that $p<i<q$ and $C_{i-1}\subset C_{p}\cup C_{q}$. From the inclusion
$C_{i}\supset(C_{i-1}\cap C_{i})\cup(C_{i}\cap C_{i+1}\cap\cdots\cap C_{q}),$
and the equalities $\big{|}\bigcap_{j=i}^{q}C_{j}\big{|}=d-(q-i)$, $\big{|}\bigcap_{j=i-1}^{q}C_{j}\big{|}=d-(q-(i-1))$, and
$|C_{i}|-|(C_{i-1}\cap C_{i})\cup(C_{i}\cap C_{i+1}\cap\cdots\cap C_{q})|=d-\Bigg{(}(d-1)+\Bigg{|}\bigcap_{j=i}^{q}C_{j}\Big{|}-\Bigg{|}\bigcap_{j=i-1}^{q}C_{j}\Bigg{|}\Bigg{)}=0,$
one obtains $C_{i}=(C_{i-1}\cap C_{i})\cup(C_{i}\cap C_{i+1}\cap\cdots\cap C_{q})\subset(C_{p}\cup C_{q})\cup C_{q}=C_{p}\cup C_{q}$. ∎
Corollary 4.7.
Let $G$ be an unmixed graph. If $\mathcal{G}$ is a tree, then $I_{c}(G)$ has a linear presentation if and only if for every path $\{C_{1},\ldots,C_{n}\}$ of $\mathcal{G}$ one has $\alpha_{0}(G)\geq n-1$ and
$\Big{|}\bigcap_{j=1}^{n}C_{j}\Big{|}=\alpha_{0}(G)-(n-1).$
Proof.
(a)$\Rightarrow$(b) Let $\{C_{1},\dots,C_{n}\}$ be a path of $\mathcal{G}$. By Proposition 3.4, there is a path $\mathcal{P}$ of $\mathcal{G}$ joining $C_{1}$ and $C_{n}$ whose vertices are in $V(\mathcal{G}^{(C_{1},C_{n})})$. As $\mathcal{G}$ is a tree, there is a unique path between $C_{1}$ and $C_{n}$. Thus, $\mathcal{P}=\{C_{1},\dots,C_{n}\}$, and consequently $C_{i}\subset C_{1}\cup C_{n}$ for any $1\leq i\leq n$. Given integers $1\leq p\leq i\leq q\leq n$, as $\mathcal{G}$ is a tree, the unique path between $C_{p}$ and $C_{q}$ is $\{C_{p},\ldots,C_{q}\}$. Again, by Proposition 3.4, one has $C_{i}\subset C_{p}\cup C_{q}$ for any $p\leq i\leq q$. Hence, by Proposition 4.6, $|\bigcap_{j=1}^{n}C_{j}|=\alpha_{0}(G)-(n-1)$ and the proof is complete.
(b)$\Rightarrow$(a) Let $C_{1}$ and $C_{n}$ be two vertices of $\mathcal{G}$. As $\mathcal{G}$ is connected, there is a path $\{C_{1},\dots,C_{n}\}$ of $\mathcal{G}$ joining $C_{1}$ and $C_{n}$. Let $p$ and $q$ be any integers such that $1\leq p\leq q\leq n$. Then, $\{C_{p},\ldots,C_{q}\}$ is a path of $\mathcal{G}$ of length $q-p$ and, by hypothesis, $|\bigcap_{j=p}^{q}C_{j}|=\alpha_{0}(G)-(q-p)$. Then, by Proposition 4.6, $C_{i}\subset C_{p}\cup C_{q}$ for $1\leq p\leq i\leq q\leq n$. Then, in particular, $\mathcal{P}$ is a path in $\mathcal{G}^{(C_{1},C_{n})}$ and, by Proposition 3.4, $I_{c}(G)$ has a linear presentation. ∎
Proposition 4.8.
Let $G$ be an unmixed graph and let $\mathcal{G}$ be the graph of the ideal of covers $I_{c}(G)$ of $G$. If $\mathcal{G}=\{C_{1},\ldots,C_{n}\}$ is a path of $\mathcal{G}$ of length $n-1$, then $I_{c}(G)$ has a linear presentation if and only if there exists a set of exchange edges $\{\{t_{k_{i}},t_{\ell_{i}}\}\}_{i=1}^{n-1}$ of $G$ such that
  1. (a)
    $C_{i}=(C_{i-1}\setminus\{t_{k_{i-1}}\})\cup\{t_{\ell_{i-1}}\}$, $t_{k_{i-1}}\in C_{i-1}$, $t_{\ell_{i-1}}\notin C_{i-1}$ for $i=2,\ldots,n$,
  2. (b)
    $|\bigcap_{j=p}^{q}C_{j}|=\alpha_{0}(G)-(q-p)$ for all $1\leq p\leq q\leq n$,
  3. (c)
    $C_{1}=\{t_{k_{i}}\}_{i=1}^{n-1}$, $C_{n}=\{t_{\ell_{i}}\}_{i=1}^{n-1}$, $V(G)=C_{1}\cup C_{n}$, and $|C_{1}\cap C_{n}|=2\alpha_{0}(G)-|V(G)|$.
Proof.
$\Rightarrow$) By Theorem 3.5, there exists a set of exchange edges of $G$ that satisfies (a). As the path $\mathcal{G}$ is a tree, by Corollary 4.7, one obtains condition (b). To prove (c) note that $\bigcap_{i=1}^{n}C_{i}=\emptyset$ because the edge ideal of $G$ is given by $I(G)=\bigcap_{i=1}^{n}(C_{i})$ and $I(G)$ contains no variables. Then, making $p=1$ and $q=n$ in (b), we get that $\alpha_{0}(G)=n-1$. Hence to show the equalities $C_{1}=\{t_{k_{i}}\}_{i=1}^{n-1}$ and $C_{n}=\{t_{\ell_{i}}\}_{i=1}^{n-1}$ it suffices to show the inclusions $C_{1}\subset\{t_{k_{i}}\}_{i=1}^{n-1}$ and $C_{n}\subset\{t_{\ell_{i}}\}_{i=1}^{n-1}$. To show the first inclusion take $t\in C_{1}$. If $t\notin\{t_{k_{i}}\}_{i=1}^{n-1}$, by part (a), one has
$t\in C_{1}\setminus\{t_{k_{1}}\}\subset C_{2},\ldots,t\in C_{n-1}\setminus\{t_{k_{n-1}}\}\subset C_{n},$
and consequently $t\in\bigcap_{i=1}^{n}C_{i}$, a contradiction. Thus, $t\in\{t_{k_{i}}\}_{i=1}^{n-1}$. To show the second inclusion take $t^{\prime}\in C_{n}$. If $t^{\prime}\notin\{t_{\ell_{i}}\}_{i=1}^{n-1}$, by part (a), one has
$t^{\prime}\in C_{n-1}\setminus\{t_{k_{n-1}}\}\subset C_{n-1},\ldots,t^{\prime}\in C_{2}\setminus\{t_{k_{2}}\}\subset C_{2},\,t^{\prime}\in C_{1}\setminus\{t_{k_{1}}\}\subset C_{1},$
and consequently $t^{\prime}\in\bigcap_{i=1}^{n}C_{i}$, a contradiction. Thus, $t^{\prime}\in\{t_{\ell_{i}}\}_{i=1}^{n-1}$. As $\mathcal{G}$ is a path, the set of exchange edges of $G$ is $\{\{t_{k_{i}},t_{\ell_{i}}\}\}_{i=1}^{n-1}$. Hence, by Proposition 4.5, the exchange edges cover $V(G)$, that is, $V(G)=C_{1}\cup C_{n}$. Therefore, $|V(G)|=2\alpha_{0}(G)-|C_{1}\cap C_{n}|$.
$\Leftarrow$) As $\mathcal{G}$ is a path and condition (b) holds, by Corollary 4.7, the ideal of covers $I_{c}(G)$ of $G$ has a linear presentation. ∎
Proposition 4.9.
Let $H_{r}$ be the graph with vertex set $V(H_{r})=\{t_{1},\ldots,t_{r},t_{r+1},\ldots,t_{2r}\}$ and edge set $E(H_{r})=\{\{t_{i},t_{r+i}\}\}_{i=1}^{r}$. The following hold.
  1. (a)
    The graph $\mathcal{H}_{r}$ of $I_{c}(H_{r})$ is a bipartite $r$-regular graph with $2^{r}$ vertices and $r2^{r-1}$ edges.
  2. (b)
    If $G$ is an unmixed bipartite graph without isolated vertices and $r=\alpha_{0}(G)$, then $G$ has a perfect matching $\{\{t_{i},t_{r+i}\}\}_{i=1}^{r}$, and the graph $\mathcal{G}$ of $I_{c}(G)$ is a subgraph of $\mathcal{H}_{r}$. In particular $\mathcal{G}$ is bipartite.
Proof.
(a) We give a recursive method to construct $\mathcal{H}_{r}$. We proceed by induction on $r\geq 2$. If $r=2$, then the minimal vertex covers of $H_{2}$ are
$C_{1}=\{t_{1},t_{2}\},\ C_{2}=\{t_{3},t_{4}\},\ C_{3}=\{t_{1},t_{4}\},\ C_{4}=\{t_{2},t_{3}\}.$
As the graph $\mathcal{H}_{2}$ is the $4$-cycle $\{C_{1},C_{3},C_{2},C_{4},C_{1}\}$. The graph $\mathcal{H}_{2}$ is bipartite with bipartition $\{C_{1},C_{2}\}$, $\{C_{3},C_{4}\}$. Assume $r>2$ and that the result holds for $\mathcal{H}_{r}$. Let $p=2^{r-1}$ and let $\{C_{1},\ldots,C_{p}\}$, $\{C_{p+1},\ldots,C_{2p}\}$ be a bipartition of the bipartite graph $\mathcal{H}_{r}$. Consider the graph $H_{r+1}$ obtained from $H_{r}$ by adding the edge $\{t_{2r+1},t_{2r+2}\}$. Then, $\alpha_{0}(H_{r+1})=r+1$. It is not hard to see that the graph $\mathcal{H}_{r+1}$ is bipartite with bipartition
$\{C_{1}\cup\{t_{2r+2}\},\ldots,\quad C_{p}\cup\{t_{2r+2}\},\ \ C_{p+1}\cup\{t_{2r+1}\},\ldots,C_{2p}\cup\{t_{2r+1}\}\},$
$\{C_{p+1}\cup\{t_{2r+2}\},\ldots,C_{2p}\cup\{t_{2r+2}\},\ C_{1}\cup\{t_{2r+1}\},\ldots,\quad C_{p}\cup\{t_{2r+1}\}\}.$
To determine the edges of $\mathcal{H}_{r+1}$ consider the following sets:
$E_{1}$ =$=\{\{C_{i}\cup\{t_{2r+2}\},C_{j}\cup\{t_{2r+2}\}\}\mid\{C_{i},C_{j}\}\in E(\mathcal{H}_{r})\},$
$E_{2}$ =$=\{\{C_{i}\cup\{t_{2r+1}\},C_{j}\cup\{t_{2r+1}\}\}\mid\{C_{i},C_{j}\}\in E(\mathcal{H}_{r})\},$
$E_{3}$ =$=\{\{C_{i}\cup\{t_{2r+2}\},C_{i}\cup\{t_{2r+1}\}\}\mid i=1,\ldots,2p\}.$
Note that $E(\mathcal{H}_{r+1})=E_{1}\cup E_{2}\cup E_{3}$ and, by induction, it follows that $\mathcal{H}_{r+1}$ is a bipartite $(r+1)$-regular graph and $E(\mathcal{H}_{r+1})=r2^{r-1}+r2^{r-1}+2^{r}=(r+1)2^{r}$.
(b) Let $C$ be any vertex of $\mathcal{G}$. Then, as $G$ is unmixed, $C$ is a vertex cover of $G$ with $r=\alpha_{0}(G)$ elements. By [32, Theorem 1.1], $G$ has a perfect matching $\{\{t_{i},t_{r+i}\}\}_{i=1}^{r}$. Then, $C$ is a vertex cover of $H_{r}$, that is, $C$ is a vertex of $\mathcal{H}_{r}$. Thus, $V(\mathcal{G})\subset V(\mathcal{H}_{r})$. Let $\{C,D\}$ be any edge of $\mathcal{G}$. Then, $|C\cap D|=\alpha_{0}(G)-1$ and, since $\alpha_{0}(G)=\alpha_{0}(H_{r})=r$, $\{C,D\}$ is an edge of $\mathcal{H}_{r}$. This proves that $\mathcal{G}$ is a subgraph of $\mathcal{H}_{r}$. Hence, by part (a), $\mathcal{G}$ is bipartite. ∎
Proposition 4.10.
Let $G$ be an unmixed graph. Then, the set $F$ of column vectors of $\mathcal{L}\mathcal{S}(I_{c}(G))$ is a minimal system of generators for $S(F)$ if and only if $\mathcal{G}=\mathcal{G}_{I_{c}(G)}$ has no strong $3$-cycles.
Proof.
Let $\{C_{1},\ldots,C_{r}\}$ be the set of minimal vertex cover of $G$. If $C_{j}=(C_{i}\setminus\{t_{k^{\prime}}\})\cup\{t_{k}\}$, we set $f^{i,j}=t_{k}e^{\prime}_{i}-t_{k^{\prime}}e^{\prime}_{j}$, where $e^{\prime}_{1},\ldots,e^{\prime}_{r}$ are the unit vectors of $S^{r}$. Thus,
$F=\{f^{i,j}\mid\{C_{i},C_{j}\}\in E(\mathcal{G})\mbox{ and }i<j\}$
is the set of column vectors of the linear syzygy matrix $\mathcal{L}\mathcal{S}(I_{c}(G))$ of $I_{c}(G)$.
$\Rightarrow$) We argue by contradiction assuming there is $C^{3}=\{C_{1},C_{2},C_{3},C_{1}\}$ a strong $3$-cycle of $\mathcal{G}$. Since $\{C_{1},C_{2}\},\{C_{1},C_{3}\}\in E(\mathcal{G})$, we can write $C_{2}=(C_{1}\setminus\{t_{j}\})\cup\{t_{i}\}$ and $C_{3}=(C_{1}\setminus\{t_{j^{\prime}}\})\cup\{t_{i^{\prime}}\}$, where $t_{j}\in C_{1}$, $t_{i}\notin C_{1}$, $t_{j^{\prime}}\in C_{1}$, $t_{i^{\prime}}\notin C_{1}$. As $C^{3}$ is strong, then $\{t_{i}\}=C_{2}\setminus C_{1}=C_{3}\setminus C_{1}=\{t_{i^{\prime}}\}$, $t_{i}=t_{i^{\prime}}$ and $t_{j}\neq t_{j^{\prime}}$. Note that $t_{j^{\prime}}\in C_{2}=(C_{1}\setminus\{t_{j}\})\cup\{t_{i}\}$ because $t_{j}\neq t_{j^{\prime}}$, and $t_{j}\notin C_{2}$ because $t_{i}\neq t_{j}$. We claim that $C_{3}=(C_{2}\setminus\{t_{j^{\prime}}\})\cup\{t_{j}\}$. As both sides have the same cardinality, it suffices to show the inclusion $\supset$”. Take $t\in(C_{2}\setminus\{t_{j^{\prime}}\})\cup\{t_{j}\}$. If $t=t_{j}$, then $t_{j}\in C_{3}$ because $t_{j}\neq t_{j^{\prime}}$ and $t_{j}\in C_{1}$. If $t=t_{i}$, then $t\in C_{3}$ because $C_{3}\setminus C_{1}=\{t_{i}\}$. If $t\in C_{2}\setminus\{t_{j^{\prime}}\}$, $t\neq t_{i}$, then $t\in C_{1}\setminus\{t_{j^{\prime}}\}\subset C_{3}$ because $C_{2}=(C_{1}\setminus\{t_{j}\})\cup\{t_{i}\}$. This proves the claim. Hence, $f^{1,2}=t_{i}e^{\prime}_{1}-t_{j}e^{\prime}_{2}$, $f^{1,3}=t_{i}e^{\prime}_{1}-t_{j^{\prime}}e^{\prime}_{3}$ and $f^{2,3}=t_{j}e^{\prime}_{2}-t_{j^{\prime}}e^{\prime}_{3}$. Thus, $f^{2,3}=f^{1,3}-f^{1,2}\in S(f^{1,2},f^{1,3})$. Therefore, $F$ is not a minimal system of generators for $S(F)$, a contradiction.
$\Leftarrow$) Arguing by contradiction we may assume $f^{1,2}=t_{q}e^{\prime}_{1}-t_{q^{\prime}}e^{\prime}_{2}\in S(F^{\prime})$, where $F^{\prime}=F\setminus\{f^{1,2}\}$. Thus, $C_{2}=(C_{1}\setminus\{t_{q^{\prime}}\})\cup\{t_{q}\}$ and $f^{1,2}=\sum_{f^{i,j}\in F^{\prime}}g_{i,j}f^{i,j}$, where $g_{i,j}\in S$ for each $f^{i,j}\in F^{\prime}$. Now, taking the inner product of $f^{1,2}$ with $e^{\prime}_{1}$, we obtain
(4.1) $t_{q}=\langle f^{1,2},e^{\prime}_{1}\rangle=\sum_{f^{1,j}\in F^{\prime}}g_{1,j}\langle f^{1,j},e^{\prime}_{1}\rangle=\sum_{f^{1,j}\in F^{\prime}}g_{1,j}t_{k_{1,j}},$
where $f^{1,j}=t_{k_{1,j}}e^{\prime}_{1}-t_{k^{\prime}_{1,j}}e^{\prime}_{i}$. If $t_{k_{1,j}}\neq t_{q}$ for each $f^{1,j}\in F^{\prime}$. Then, making $t_{k_{1,j}}=0$ for all $j\neq 1$ and $t_{q}=1$ in Eq. (4.1), we obtain $1=0$, a contradiction. Then, there is $f^{1,j}\in F^{\prime}$ such that $t_{k_{1,j}}=t_{q}$. Without loss of generality, we may assume $j=3$, then $f^{1,3}=t_{q}e^{\prime}_{1}-t_{k^{\prime}_{1,3}}e^{\prime}_{3}$. Consequently, $C_{3}=(C_{1}\setminus\{t_{k^{\prime}_{1,3}}\})\cup\{t_{q}\}$. We can proceed as in the first part of the proof to show the equality $C_{3}=(C_{2}\setminus\{t_{k^{\prime}_{1,3}}\})\cup\{t_{q}\}$. Then, $C_{2}\setminus C_{1}=\{t_{q}\}=C_{3}\setminus C_{1}$, and $\{C_{1},C_{2},C_{3},C_{1}\}$ is a strong $3$-cycle of $\mathcal{G}$, a contradiction. Therefore, the set of columns of $\mathcal{L}\mathcal{S}(I_{c}(G))$ is a minimal system of generators for $S(F)$. ∎
5. Linearly presented ideals of covers
Let $G$ be an unmixed graph. In this section we show that the ideal of covers $I_{c}(G)$ of $G$ is linearly presented if $G$ has no induced 4-cycles and classify combinatorially when $I_{c}(G)$ is linearly presented for graphs without $3$- and $5$-cycles.
Proposition 5.1.
An edge $e$ has the (P) property if and only if $|e\cap C|=1$ for each minimal vertex cover $C$ of $G$.
Proof.
$\Rightarrow$) By contradiction suppose there is a minimal vertex cover $C$ of $G$ such that $e\subset C$. We assume $e=\{t_{1},t_{2}\}$, then there are $t_{1}^{\prime},t_{2}^{\prime}\in V(G)\setminus C$ such that $\{t_{1}^{\prime},t_{1}\},\{t_{2}^{\prime},t_{2}\}\in E(G)$, since $C$ is minimal. Furthermore, as $e$ has the (P) property, then $\{t_{1}^{\prime},t_{2}^{\prime}\}\in E(G)$. Hence, $\{t_{1}^{\prime},t_{2}^{\prime}\}\cap C\neq\emptyset$, since $C$ is a vertex cover, a contradiction, since $t_{1}^{\prime},t_{2}^{\prime}\in V(G)\setminus C$.
$\Leftarrow$) We let $e=\{t_{1},t_{2}\}$. To show that $e$ has the (P) property let $\{t_{1},t^{\prime}_{1}\}$ and $\{t_{2},t^{\prime}_{2}\}$ be two edges of $G$ distinct from $e$. We argue by contradiction assuming $\{t^{\prime}_{1},t^{\prime}_{2}\}$ is not an edge of $G$. Then, there exists a minimal vertex cover $C$ of $G$ such that $C\cap\{t_{1}^{\prime},t_{2}^{\prime}\}=\emptyset$. Then, $t_{i}\in C$ for $i=1,2$ and $|C\cap e|=2$, a contradiction. ∎
Proposition 5.2.
If $C^{4}$ is a $4$-cycle of a graph $G$, then the following conditions are equivalent:
  1. (a)
    Each edge of $C^{4}$ has the (P) property.
  2. (b)
    $C^{4}$ has two disjoint edges with the (P) property.
  3. (c)
    $|V(C^{4})\cap C|=2$ for every minimal vertex cover $C$ of $G$.
Proof.
(a)$\Rightarrow$(b) This implication is clear.
(b)$\Rightarrow$(c) There are $\epsilon_{1},\epsilon_{2}\in E(C^{4})$ such that they have the (P) property and $\epsilon_{1}\cap\epsilon_{2}=\emptyset$. Then, $\epsilon_{1}\cup\epsilon_{2}=V(C^{4})$. We take a minimal vertex cover $C$ of $G$. Thus, by Proposition 5.1, $|\epsilon_{i}\cap C|=1$ for $i=1,2$. Hence, $|V(C^{4})\cap C|=2$, since $\epsilon_{1}\cap\epsilon_{2}=\emptyset$.
(c)$\Rightarrow$(a) We argue by contradiction assuming that $C^{4}=\{t_{1},t_{2},t_{3},t_{4},t_{1}\}$ is a $4$-cycle of $G$ such that $\{t_{1},t_{2}\}$ does not have the (P) property. Then, there are $\{t_{1},t^{\prime}_{1}\},\{t_{2},t^{\prime}_{2}\}\in E(G)$ such that $\{t^{\prime}_{1},t^{\prime}_{2}\}\notin E(G)$. Consequently, there is a maximal stable set $A$ of $G$ such that $\{t^{\prime}_{1},t^{\prime}_{2}\}\subset A$. So, $t_{1}\notin A$ and $t_{2}\notin A$. Furthermore $t_{3}\notin A$ or $t_{4}\notin A$, since $\{t_{3},t_{4}\}\in E(G)$. We may assume $t_{3}\notin A$. Since $A$ is a maximal stable set, we have that $C:=V(G)\setminus A$ is a minimal vertex cover and $t_{1},t_{2},t_{3}\in C$. A contradiction, since $|V(C^{4})\cap C|=2$. ∎
Lemma 5.3.
If $C^{4}=\{t_{1},t_{2},t_{3},t_{4},t_{1}\}$ is a $4$-cycle of a graph $G$ and each edge of $C^{4}$ has the (P) property, then $N_{G}(t_{1})=N_{G}(t_{3})$, $N_{G}(t_{2})=N_{G}(t_{4})$, and $C^{4}$ is an induced cycle.
Proof.
To show the inclusion $N_{G}(t_{1})\subset N_{G}(t_{3})$ take $t\in N_{G}(t_{1})$, then $\{t,t_{1}\}$ and $\{t_{2},t_{3}\}$ are edges of $G$. We may assume $t\neq t_{2}$, otherwise $t\in N_{G}(t_{3})$. As $\{t_{1},t_{2}\}$ has the (P) property, $t\neq t_{3}$ and $\{t,t_{3}\}\in E(G)$. Thus, $t\in N_{G}(t_{3})$. To show the inclusion $\supset$” take $t^{\prime}\in N_{G}(t_{3})$, then $\{t^{\prime},t_{3}\}$ and $\{t_{1},t_{4}\}$ are edges of $G$. We may assume $t^{\prime}\neq t_{4}$, otherwise $t^{\prime}\in N_{G}(t_{1})$. As $\{t_{3},t_{4}\}$ has the (P) property, $t^{\prime}\neq t_{1}$ and $\{t^{\prime},t_{1}\}\in E(G)$. Thus, $t^{\prime}\in N_{G}(t_{1})$. The equality $N_{G}(t_{2})=N_{G}(t_{4})$ follows using similar arguments. From the two equalities, we get that $C^{4}$ is an induced cycle since $G$ is a simple graph. ∎
Theorem 5.4.
Let $G$ be an unmixed graph without isolated vertices and let $I_{c}(G)$ be its ideal of covers. If the graph $\mathcal{G}_{I_{c}(G)}$ of $I_{c}(G)$ is connected, then
  1. (a)
    $G$ has no duplicated vertices, and
  2. (b)
    Every $4$-cycle of $G$ has an edge that does not satisfy the (P) property.
Proof.
(a) Let $E(G^{\vee})$ be the set of the minimal vertex covers of $G$. By contradiction suppose $t$ and $t^{\prime}$ are two duplicated vertices of $G$. We let
$A=\{C\in E(G^{\vee})\mid t,t^{\prime}\in C\}\mbox{ and }B=\{C\in E(G^{\vee})\mid t\notin C\mbox{ and }t^{\prime}\notin C\}.$
Now assume there is $C\in E(G^{\vee})\setminus(A\cup B)$. Thus, |$|\{t,t^{\prime}\}\cap C|=1$. We may assume $t\in C$ and $t^{\prime}\notin C$, then $N_{G}(t^{\prime})\subset C$. Since $t$ and $t^{\prime}$ are duplicated, $N_{G}(t)=N_{G}(t^{\prime})$. Consequently, $C\setminus\{t\}$ is a vertex cover. A contradiction, since $C$ is minimal. Hence, $E(G^{\vee})=A\cup B$. As $t$ is not an isolated vertex, by Lemma 4.4, there are minimal vertex covers $C$ and $C^{\prime}$ such that $t\in C^{\prime}$ and $t\notin C$. Then, $C^{\prime}\notin B$ (resp. $C\notin A$) because $t\in C^{\prime}$ (resp. $t\notin C$). Thus, by the equality $E(G^{\vee})=A\cup B$, one has $C^{\prime}\in A$ and $C\in B$. As $\mathcal{G}_{I_{c}(G)}$ is connected, there is a path $\mathcal{P}=\{C_{1}=C^{\prime},C_{2},\ldots,C_{n}=C\}$ between $C^{\prime}$ and $C$. Pick $i$ such that $C_{i}\in A$ and $C_{i+1}\in B$. Hence, $\{t,t^{\prime}\}\subset C_{i}\setminus C_{i+1}$ and $|C_{i}\setminus C_{i+1}|\geq 2$, a contradiction since $\{C_{i},C_{i+1}\}\in E(\mathcal{G}_{I_{c}(G)})$.
(b) By contradiction assume that $C^{4}$ is a 4-cycle of $G$ whose edges have the (P) property, then $G$ has duplicated vertices by Lemma 5.3, a contradiction to part (a). ∎
Remark 5.5.
If $C^{4}=\{t_{1},t_{2},t_{3},t_{4},t_{1}\}$ is a $4$-cycle with chord $\{t_{1},t_{3}\}$, then $\{t_{1},t_{2}\}$ does not satisfy the (P) property, since $\{t_{1},t_{3}\},\{t_{2},t_{3}\}\in E(G)$, $\{t_{3},t_{3}\}\notin E(G)$ and recalling that $G$ is a simple graph. Therefore, condition (b) of Theorem 5.4 is equivalent to (b’) Every induced $4$-cycle has an edge without the (P) property.
Proposition 5.6.
If $P$ is a perfect matching with the (P) property of a graph $G$, then $P$ is a perfect matching of $G$ of König type.
Proof.
We assume $P=\{\epsilon_{1},\ldots,\epsilon_{m}\}$ and we take $C$ a minimal vertex cover of $G$. Then, by Proposition 5.1, $|\epsilon_{i}\cap C|=1$ for each $1\leq i\leq m$. Thus, $|C|=m$, since $P$ is a partition of $V(G)$. Hence, $G$ is unmixed and $\alpha_{0}(G)=m=|P|$, that is, $P$ is a perfect matching of König type. ∎
Remark 5.7.
By Proposition 5.6, we can remove the hypothesis that $P$ is of König type in parts (a) and (b) of Proposition 2.7.
Theorem 5.8.
Let $G$ be an unmixed König graph without isolated vertices. The following conditions are equivalent.
  1. (a)
    G is Cohen–Macaulay.
  2. (b)
    $I_{c}(G)$ is linearly presented.
  3. (c)
    $\mathcal{G}_{I_{c}(G)}$ is connected.
  4. (d)
    $G$ has no duplicated vertices.
  5. (e)
    Every induced $4$-cycle of $G$ has an edge without the (P) property.
  6. (f)
    $G$ has a unique perfect matching.
Proof.
(a) $\Rightarrow$ (b) By Eagon-Reiner criterion $I_{c}(G)$ has a linear resolution (Theorem 2.4). Hence, $I_{c}(G)$ is linearly presented.
(b) $\Rightarrow$ (c) $\Rightarrow$ (d) $\Rightarrow$ (e) These implications follow from Proposition 3.4, Theorem 5.4, and Lemma 5.3, respectively.
(e) $\Rightarrow$ (a) By Proposition 2.7(a), $G$ has a perfect matching $P$ of König type with the (P) property. By contradiction assume $G$ is not Cohen–Macaulay. Then, by Proposition 2.7(b), there is a $4$-cycle $C^{4}$ with two edges in $P$. Thus, by Proposition 5.2 (cf. Lemma 5.3), each edge of $C^{4}$ has the (P) property. Hence, by Remark 5.5, $C^{4}$ is induced. A contradiction to (e).
(a)$\Leftrightarrow$(f) An unmixed König graph $G$ without isolated vertices has a perfect matching of König type [22, Lemma 2.3]. Thus, $G$ is very well-covered. Hence, by [7, Theorem 0.2], $G$ is Cohen–Macaulay if and only if $G$ has a unique perfect matching. ∎
Lemma 5.9.
Let $G,H$ be unmixed graphs such that $H\subset G$. If $V(H)=V(G)$ and $\alpha_{0}(G)=\alpha_{0}(H)$, then $\mathcal{G}_{I_{c}(G)}\subset\mathcal{G}_{I_{c}(H)}$.
Proof.
We set $\mathcal{G}_{1}=\mathcal{G}_{I_{c}(G)}$ and $\mathcal{G}_{2}=\mathcal{G}_{I_{c}(H)}$. We take $C\in V(\mathcal{G}_{1})$, then $C$ is a minimal vertex cover of the graph $G$. So, $C\subset V(G)=V(H)$ and $|C|=\alpha_{0}(G)=\alpha_{0}(H)$. Thus, $C$ is a minimal vertex cover of $H$, since $E(H)\subset E(G)$. Hence, $C\in V(\mathcal{G}_{2})$. Furthermore, if $\{C_{1},C_{2}\}\in E(\mathcal{G}_{1})$, then $\{C_{1},C_{2}\}\in E(\mathcal{G}_{2})$. Therefore, as $H$ is unmixed, $\mathcal{G}_{1}\subset\mathcal{G}_{2}$. ∎
Theorem 5.10.
If $G$ is a Cohen–Macaulay König graph without isolated vertices, then
  1. (a)
    $G$ has a unique perfect matching $P$ which is the set of exchange edges of $G$, and
  2. (b)
    $\mathcal{G}_{I_{c}(G)}$ is a bipartite graph.
Proof.
(a) By induction on $|V(G)|$. From Proposition 2.7(b), $G$ has a perfect matching $P$ with the (P) property. By Proposition 2.10, there is $t\in V(G)$ such that $N_{G}(t)=\{t^{\prime}\}$. So, $\epsilon_{1}=\{t,t^{\prime}\}\in P$, since $P$ is a partition of $V(G)$. Now, we let $G^{\prime}=G\setminus N_{G}[t]$, then $G^{\prime}$ is Cohen–Macaulay because being Cohen–Macaulay is closed under c-minors [33, 6.3.53]. Furthermore, $P^{\prime}:=P\setminus\{\epsilon_{1}\}$ is a perfect matching of $G^{\prime}$ with the (P) property in $G^{\prime}$, since $G^{\prime}=G\setminus\{t,t^{\prime}\}$. Thus, by Proposition 5.1, $\alpha_{0}(G^{\prime})=|P^{\prime}|$, since $P^{\prime}$ is a partition of $V(G^{\prime})$. Then, $G^{\prime}$ is a König graph. Hence, by induction hypothesis, $P^{\prime}$ is the unique perfect matching of $G^{\prime}$ and $P^{\prime}=Ex(G^{\prime})$, where $Ex(G)$ and $Ex(G^{\prime})$ are the exchange edges of $G$ and $G^{\prime}$, respectively.
We let $\mathcal{C}=E(G^{\vee})$ and $\mathcal{C}^{\prime}=E((G^{\prime})^{\vee})$ be the set of minimal vertex covers of the graphs $G$ and $G^{\prime}$, respectively. Take $\{t_{1},t_{2}\}\in P^{\prime}=Ex(G^{\prime})$, then there are $C^{\prime}_{1},C^{\prime}_{2}\in\mathcal{C}^{\prime}$ such that $C^{\prime}_{2}=(C^{\prime}_{1}\setminus\{t_{1}\})\cup\{t_{2}\}$. Consequently, $C_{1}=C^{\prime}_{1}\cup\{t^{\prime}\},C_{2}=C^{\prime}_{2}\cup\{t^{\prime}\}\in\mathcal{C}$, since $G^{\prime}=G\setminus\{t,t^{\prime}\}$ and $N_{G}(t)=\{t^{\prime}\}$. Furthermore, $C_{2}=(C_{1}\setminus\{t_{1}\})\cup\{t_{2}\}$, then $\{t_{1},t_{2}\}\in Ex(G)$. This implies, $P^{\prime}\subset Ex(G)$. On the other hand, by Lemma 4.4, there is $\overline{C}\in\mathcal{C}$ such that $t\in\overline{C}$, then $(\overline{C}\setminus\{t\})\cup\{t^{\prime}\}\in\mathcal{C}$, since $N_{G}(t)=\{t^{\prime}\}$. Hence, $\epsilon_{1}=\{t,t^{\prime}\}\in Ex(G)$ and $P\subset Ex(G)$.
Now, we will prove that $P=Ex(G)$. By contradiction assume there is $\{\overline{t}_{1},\overline{t}_{2}\}\in Ex(G)\setminus P$, then there are $C_{1},C_{2}\in\mathcal{C}$ such that $C_{2}=(C_{1}\setminus\{\overline{t}_{1}\})\cup\{\overline{t}_{2}\}$. By Proposition 5.1,
$|C_{1}\cap\epsilon_{1}|=1=|C_{2}\cap\epsilon_{1}|,$
since $\epsilon_{1}$ has the (P) property. Thus, $C_{1}=C^{\prime}_{1}\cup\{\tilde{t}\}$ and $C_{2}=C^{\prime}_{2}\cup\{\tilde{t}^{\prime}\}$, where $C^{\prime}_{1},C^{\prime}_{2}\subset E(G^{\prime})$ and $\{\tilde{t},\tilde{t}^{\prime}\}\subset\{t,t^{\prime}\}$. So, $C^{\prime}_{1},C^{\prime}_{2}\in\mathcal{C}^{\prime}$, since $C_{1},C_{2}\in\mathcal{C}$ and $\alpha_{0}(G^{\prime})=|P^{\prime}|=|P|-1=\alpha_{0}(G)-1$. If $\tilde{t}\neq\tilde{t}^{\prime}$, then $C^{\prime}_{1}=C^{\prime}_{2}$ and $\{\overline{t}_{1},\overline{t}_{2}\}=\{t,t^{\prime}\}=\epsilon_{1}$. A contradiction, since $\{\overline{t}_{1},\overline{t}_{2}\}\notin P$. Consequently, $\tilde{t}=\tilde{t}^{\prime}$ and $C^{\prime}_{2}=(C^{\prime}_{1}\setminus\{\overline{t}_{1}\})\cup\{\overline{t}_{2}\}$. Hence, $\{\overline{t}_{1},\overline{t}_{2}\}\in Ex(G^{\prime})=P^{\prime}\subset P$. This is a contradiction. Therefore, $P=Ex(G)$.
(b) By Proposition 5.1, $\alpha_{0}(G)=|P|=\alpha_{0}(P)$. Thus, by Lemma 5.9, $\mathcal{G}_{I_{c}(G)}\subset\mathcal{G}_{I_{c}(P)}$, since $P\subset G$ and $V(P)=V(G)$. Furthermore, by Proposition 4.9, the graph $\mathcal{G}_{I_{c}(P)}$ is bipartite. Hence, $\mathcal{G}_{I_{c}(G)}$ is bipartite. ∎
Recall that if $G$ is a graph and $L\subset V(G)$, then $G[L]$ denotes the subgraph of $G$ induced by $L$. The edge set of $G[L]$ is $L$ and $e\in E(G[L])$ if and only if $e\subset L$ and $e\in E(G)$.
Lemma 5.11.
Let $G$ be an unmixed graph with distinct minimal vertex covers $C_{1}$ and $C_{2}$. If $C^{\prime}$ is a vertex cover of $G^{\prime}:=G[C_{1}\triangle C_{2}]$, then the following conditions hold:
  1. (a)
    $C^{\prime}\cup A$ is a vertex cover of $G$, where $A:=C_{1}\cap C_{2}$.
  2. (b)
    $G^{\prime}$ is a bipartite graph with a perfect matching $P$ and bipartition $(B_{1},B_{2})$ such that $|B_{1}|=|B_{2}|=|P|=\alpha_{0}(G^{\prime})$, where $B_{1}:=C_{1}\setminus C_{2}$ and $B_{2}:=C_{2}\setminus C_{1}$.
Proof.
(a) Let $C^{\prime}$ be any vertex cover of $G^{\prime}$. We prove that $C^{\prime}\cup A$ is a vertex cover of $G$, where $A=C_{1}\cap C_{2}$. To show this, take $\epsilon\in E(G)$. We may assume $\epsilon\cap A=\emptyset$. Since $C_{i}$ is vertex cover, there is $t_{i}\in\epsilon\cap C_{i}$ for $i=1,2$. Then, $t_{1}\notin C_{2}$ and $t_{2}\notin C_{1}$, since $\epsilon\cap(C_{1}\cap C_{2})=\epsilon\cap A=\emptyset$. So, $t_{1}\in C_{1}\setminus C_{2}$ and $t_{2}\in C_{2}\setminus C_{1}$. Thus, $\epsilon=\{t_{1},t_{2}\}\subset C_{1}\triangle C_{2}=V(G^{\prime})$. Hence, $\epsilon\cap C^{\prime}\neq\emptyset$, since $C^{\prime}$ is a vertex cover of $G^{\prime}$. Therefore, $C^{\prime}\cup A$ is a vertex cover of $G$.
(b) We have $B_{1}=C_{1}\setminus C_{2}$ and $B_{2}=C_{2}\setminus C_{1}$. In particular $B_{1}\cap C_{2}=\emptyset$ and $B_{2}\cap C_{1}=\emptyset$, then $B_{1}$ and $B_{2}$ are stable sets of $G$, since $C_{2}$ and $C_{1}$ are vertex covers. Consequently, $G^{\prime}$ is bipartite with bipartition $(B_{1},B_{2})$, since $V(G^{\prime})=C_{1}\triangle C_{2}=B_{1}\cup B_{2}$ and $B_{1}\cap B_{2}=\emptyset$. Thus, by König’s theorem (Theorem 2.6), $G^{\prime}$ is a König graph since $G^{\prime}$ is bipartite. So, there is a matching $P$ of $G^{\prime}$ such that $|P|=\alpha_{0}(G^{\prime})$. Furthermore, by (a), we have $C^{\prime}\cup A$ is a vertex cover of $G$ , for each vertex cover $C^{\prime}$ of $G^{\prime}$. Then,
(5.1) $|C^{\prime}\cup A|\geq\alpha_{0}(G)=|C_{i}|=|B_{i}\cup A|\mbox{ for }i=1,2,$
since $C_{i}$ is a minimal vertex cover and $C_{i}=B_{i}\cup A$. Notice that $C^{\prime}\cap A\subset V(G^{\prime})\cap A=\emptyset$ and $B_{i}\cap A=\emptyset$. Consequently, by Eq. (5.1), we have $|C^{\prime}|\geq|B_{i}|$ for each vertex cover $C^{\prime}$ of $G^{\prime}$. Thus, $\alpha_{0}(G^{\prime})\geq|B_{i}|$. Also, $B_{1}$, $B_{2}$ are vertex covers of $G^{\prime}$, since $B_{2}=V(G^{\prime})\setminus B_{1}$, and $B_{2}=V(G^{\prime})\setminus B_{2}$. Hence, $|B_{1}|=\alpha_{0}(G^{\prime})=|B_{2}|$. Recall $\alpha_{0}(G^{\prime})=|P|$, then
$|V(G^{\prime})|=|B_{1}|+|B_{2}|=2\alpha_{0}(G^{\prime})=2|P|.$
Therefore, $P$ is a perfect matching of $G^{\prime}$. ∎
Theorem 5.12.
If $G$ is an unmixed graph and $J=I_{c}(G)$ is not linearly presented, then there are distinct minimal vertex covers $C_{1}$ and $C_{2}$ such that $G^{\prime}=G[C_{1}\triangle C_{2}]$ is an unmixed bipartite graph without isolated vertices and $G^{\prime}$ is not Cohen-Macaulay.
Proof.
By Proposition 3.4(b), there are $C_{1},C_{2}\in E(G^{\vee})$ such that there is no path in $\mathcal{G}_{J}^{(C_{1},C_{2})}$ connecting $C_{1}$ and $C_{2}$, where $J=I_{c}(G)$. We choose $C_{1}$ and $C_{2}$ such that $|C_{1}\cap C_{2}|$ is maximal with respect to this property. One has $\alpha_{0}(G)=|C_{1}|=|C_{2}|$, since $G$ is an unmixed graph. By Lemma 5.11(b), $G^{\prime}=G[C_{1}\triangle C_{2}]$ is bipartite with a perfect matching $P$ and bipartition $(B_{1},B_{2})$ such that $|B_{1}|=|B_{2}|=|P|=\alpha_{0}(G^{\prime})$, where $B_{1}=C_{1}\setminus C_{2}$ and $B_{2}=C_{2}\setminus C_{1}$.
Now we prove that $P$ has the (P) property in $G^{\prime}$. By contradiction, suppose there is an edge $\{t,t^{\prime}\}$ in $P$ and $t_{1},t_{2}\in V(G^{\prime})$ such that $\{t,t_{1}\},\{t^{\prime},t_{2}\}\in E(G^{\prime})$ and $\{t_{1},t_{2}\}\notin E(G^{\prime})$. Since $G^{\prime}$ is bipartite, we have |$|\{t,t^{\prime},t_{1},t_{2}\}|=4$. Also, there is a maximal stable set $U^{\prime}$ of $G^{\prime}$ such that $\{t_{1},t_{2}\}\subset U^{\prime}$. Thus, $C^{\prime}=V(G^{\prime})\setminus U^{\prime}$ is a minimal vertex cover of $G^{\prime}$, $t_{1},t_{2}\notin C^{\prime}$, and $\{t,t^{\prime}\}\subset C^{\prime}$. We may assume $t\in B_{1}$ and $t^{\prime}\in B_{2}$, since $(B_{1},B_{2})$ is a bipartition of $G^{\prime}$. Then, $t_{1}\in B_{2}$ and $t_{2}\in B_{1}$. We consider the following subsets of $P$:
$P_{1}=\{\epsilon\in P\mid\mbox{ }\epsilon\subset C^{\prime}\},\mbox{ }P_{2}=\{\epsilon\in P\mid\mbox{ }\epsilon\cap C^{\prime}\subset B_{1}\}\mbox{ and }P_{3}=\{\epsilon\in P\mid\mbox{ }\epsilon\cap C^{\prime}\subset B_{2}\},$
then $P_{1},P_{2},P_{3}$ are non-empty (see the discussion below) and form a partition of $P$, since $C^{\prime}$ is a vertex cover of $G^{\prime}$. We also consider the sets
$B^{i}_{j}=V(P_{i})\cap B_{j},\mbox{ for }i=1,2,3\mbox{ and }j=1,2.$
The $B^{i}_{j}$’s are mutually disjoint, since $P_{1},P_{2},P_{3}$ is a partition of $P$ and $B_{1}\cap B_{2}=\emptyset$. Note that $|B_{1}^{i}|=|P_{i}|=|B_{2}^{i}|$ for $i=1,2,3$. Since $P$ is a perfect matching of $G^{\prime}$, there are $\{t_{1},t^{\prime}_{1}\},\{t_{2},t^{\prime}_{2}\}\in P$. Consequently, $t^{\prime}_{1}\in B_{1}\cap C^{\prime}$ and $t^{\prime}_{2}\in B_{2}\cap C^{\prime}$, since $t_{1}\in B_{2}$, $t_{2}\in B_{1}$ and $t_{1},t_{2}\notin C^{\prime}$. This implies $t^{\prime}_{1}\in B^{2}_{1}$, $t^{\prime}_{2}\in B^{3}_{2}$, $\{t_{1},t_{1}^{\prime}\}\in P_{2}$ and $\{t_{2},t_{2}^{\prime}\}\in P_{3}$. Next we prove the equality
(5.2) $C^{\prime}=B_{1}^{1}\cup B_{1}^{2}\cup B_{2}^{1}\cup B_{2}^{3}.$
The inclusion $\supset$” is clear by construction of the $P_{i}$’s. To show the inclusion $\subset$” take $t_{i}\in C^{\prime}$. There is $t_{i}^{\prime}$ such that $\{t_{i},t_{i}^{\prime}\}\in P$. If $t_{i}^{\prime}\in C^{\prime}$, then $\{t_{i},t_{i}^{\prime}\}\in P_{1}$ and $t_{i}\in B_{1}^{1}$ or $t_{i}\in B_{2}^{1}$. If $t_{i}^{\prime}\notin C^{\prime}$ and $\{t_{i},t_{i}^{\prime}\}\in P_{2}$, then $t_{i}\in B_{1}^{2}$. If $t_{i}^{\prime}\notin C^{\prime}$ and $\{t_{i},t_{i}^{\prime}\}\in P_{3}$, then $t_{i}\in B_{2}^{3}$. This proves the equality. Now, by Lemma 5.11(a), we have $C^{\prime}\cup A$ is a vertex cover of $G$ where $A:=C_{1}\cap C_{2}$. Thus, there is $C_{3}\in E(G^{\vee})$ such that $C_{3}\subset C^{\prime}\cup A$. We claim that $C^{\prime}\subset C_{3}$. We argue by contradiction assuming there is $t_{i}\in C^{\prime}\setminus C_{3}$. As $C^{\prime}$ is a minimal vertex cover of $G^{\prime}$, there is $\epsilon\in E(G^{\prime})$ such that $\epsilon\cap C^{\prime}=\{t_{i}\}$. Since $\epsilon\in E(G^{\prime})$, we have $\epsilon\cap A=\emptyset$. So, $\epsilon\cap C_{3}\subset\epsilon\cap(C^{\prime}\cup A)\subset\epsilon\cap C^{\prime}=\{t_{i}\}$. But $t_{i}\notin C_{3}$, then $\epsilon\cap C_{3}=\emptyset$. A contradiction since $C_{3}\in E(G^{\vee})$. From the claim, we can write $C_{3}=A_{1}\cup C^{\prime}$ for some $A_{1}\subset A$. Then, by Eq. (5.2) and noticing that $B_{1}=B^{1}_{1}\cup B^{2}_{1}\cup B^{3}_{1}$, we obtain that $C_{3}$ and $C_{1}$ can be partitioned as
(5.3) $C_{3}=A_{1}\cup C^{\prime}=A_{1}\cup B_{1}^{1}\cup B_{1}^{2}\cup B_{2}^{1}\cup B_{2}^{3},$
(5.4) $C_{1}=A\cup B_{1}=A_{1}\cup A_{2}\cup B^{1}_{1}\cup B^{2}_{1}\cup B^{3}_{1},\mbox{ where }A_{2}=A\setminus A_{1},$
respectively. Hence, noticing that $|C_{1}|=|C_{3}|$ and $|B^{3}_{1}|=|P_{3}|=|B^{3}_{2}|$, we obtain $|A_{2}|=|B^{1}_{2}|$. From Eqs. (5.3) and (5.4), we have $C_{1}\cap C_{3}=A_{1}\cup B^{1}_{1}\cup B^{2}_{1}$. Similarly, we obtain the equality $C_{2}\cap C_{3}=A_{1}\cup B^{1}_{2}\cup B^{3}_{2}$, since $C_{2}=A\cup B_{2}=A_{1}\cup A_{2}\cup B^{1}_{2}\cup B^{2}_{2}\cup B^{3}_{2}$. Recall $|B_{1}^{1}|=|P_{1}|=|B_{2}^{1}|$ and $|B^{1}_{2}|=|A_{2}|$, then $|C_{1}\cap C_{3}|=|A_{1}|+|B^{1}_{1}|+|B^{2}_{1}|=|A|+|B^{2}_{1}|$, since $A=A_{1}\cup A_{2}$. Similarly, we obtain that $|C_{2}\cap C_{3}|=|A|+|B^{3}_{2}|$, since $|B^{1}_{2}|=|A_{2}|$ and $A=A_{1}\cup A_{2}$. Then,
$|C_{i}\cap C_{3}|>|A|\mbox{ for }i=1,2,\mbox{ since }t^{\prime}_{1}\in B_{1}^{2}\mbox{ and }t^{\prime}_{2}\in B_{2}^{3}.$
But $C_{1}$, $C_{2}$ were chosen so that $|C_{1}\cap C_{2}|=|A|$ is maximal with respect to the property that there is no path in $\mathcal{G}_{J}^{(C_{1},C_{2})}$ connecting $C_{1}$ and $C_{2}$. Consequently, there is a path $\mathcal{P}_{i}$ between $C_{i}$ and $C_{3}$ in $\mathcal{G}_{J}^{(C_{i},C_{3})}$ for $i=1,2$. Recall that $A_{1}\subset A=C_{1}\cap C_{2}$ and $C^{\prime}\subset B_{1}\cup B_{2}\subset C_{1}\cup C_{2}$. Thus, $C_{3}\subset C_{1}\cup C_{2}$, and one has
$\mathcal{P}_{i}\subset\mathcal{G}_{J}^{(C_{i},C_{3})}\subset\mathcal{G}_{J}^{(C_{1},C_{2})}\mbox{ for }i=1,2.$
This implies, $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ is a path between $C_{1}$ and $C_{2}$ in $\mathcal{G}_{J}^{(C_{1},C_{2})}$, a contradiction. Therefore, $P$ has the (P) property. On the other hand, by König’s theorem (Theorem 2.6), $G^{\prime}$ is a König graph, since $G^{\prime}$ is bipartite. Also, $G^{\prime}$ does not have isolated vertices, since $P$ is a perfect matching of $G^{\prime}$. Hence, by Proposition 2.7(a), $G^{\prime}$ is an unmixed graph.
Now, suppose the graph $G^{\prime}$ is Cohen-Macaulay. By Proposition 2.10, there is $\overline{t}\in V(G^{\prime})$ with $\deg_{G^{\prime}}(\overline{t})=1$. We can assume $\overline{t}\in B_{1}$. Since $P$ is a perfect matching, there is $\{\overline{t},\overline{t}^{\prime}\}\in P$ with $\overline{t}^{\prime}\in B_{2}$. Consequently, $\overline{t}\neq\overline{t}^{\prime}$ and $C_{4}=(C_{1}\setminus\{\overline{t}\})\cup\{\overline{t}^{\prime}\}\in E(G^{\vee})$, since $N_{G^{\prime}}(\overline{t})=\{\overline{t}^{\prime}\}\subset C_{4}$ and $|C_{4}|=|C_{1}|$. Also, $C_{4}\subset C_{1}\cup C_{2}$ since $\overline{t}^{\prime}\in B_{2}\subset C_{2}$. Thus, $\overline{\epsilon}=\{C_{1},C_{4}\}\in E(\mathcal{G}_{J}^{(C_{1},C_{2})})$. Furthermore, $C_{4}\cap C_{2}=A\cup\{\overline{t}^{\prime}\}$ and $\overline{t}^{\prime}\notin A$, since $\overline{t}\in B_{1}=C_{1}\setminus C_{2}$ and $\overline{t}^{\prime}\in B_{2}=C_{2}\setminus A$. Then, $|C_{4}\cap C_{2}|>|A|=|C_{1}\cap C_{2}|$. This implies, because of the way we chose $C_{1}$ and $C_{2}$, that there is a path $\mathcal{P}$ between $C_{2}$ and $C_{4}$ in $\mathcal{G}_{J}^{(C_{2},C_{4})}$. So, $\mathcal{P}\in\mathcal{G}_{J}^{(C_{1},C_{2})}$, since $C_{4}\subset C_{1}\cup C_{2}$. Hence, $\mathcal{P}\cup\{\overline{\epsilon}\}$ is a path between $C_{1}$ and $C_{2}$ in $\mathcal{G}_{J}^{(C_{1},C_{2})}$, a contradiction. Therefore, $G^{\prime}$ is not Cohen-Macaulay. ∎
Theorem 5.13.
If $G$ is an unmixed graph without induced $4$-cycles, then $I_{c}(G)$ is linearly presented.
Proof.
By contradiction suppose that $I_{c}(G)$ is not linearly presented. Then, by Theorem 5.12, there are $C_{1},C_{2}\in E(G^{\vee})$ such that $G^{\prime}=G[C_{1}\triangle C_{2}]$ is not Cohen-Macaulay and $G^{\prime}$ is an unmixed bipartite graph without isolated vertices. So, by König’s theorem (Theorem 2.6), $G^{\prime}$ is a König graph. Hence, by Theorem 5.8(a)-(e), $G^{\prime}$ is Cohen-Macaulay, since $G$ does not have induced $4$-cycles. This is a contradiction. Therefore $I_{c}(G)$ is linearly presented. ∎
Corollary 5.14.
If $G$ is an unmixed graph without induced $4$-cycles, then ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1$.
Theorem 5.15.
Let $G$ be an unmixed graph such that $I_{c}(G)$ is linearly presented. If $G[C_{1}\triangle C_{2}]$ is unmixed and not Cohen-Macaulay, where $C_{1}$ and $C_{2}$ are distinct minimal vertex covers of $G$, then there is $C\in E(G^{\vee})$ such that $C_{1}\cap C_{2}\not\subset C\subset C_{1}\cup C_{2}$.
Proof.
Assume $G^{\prime}:=G[C_{1}\triangle C_{2}]$ is unmixed and not Cohen-Macaulay, where $C_{1},C_{2}\in E(G^{\vee})$ and $C_{1}\neq C_{2}$. We will show there is $C\in E(G^{\vee})$ such that $A\not\subset C\subset C_{1}\cup C_{2}$, where $A:=C_{1}\cap C_{2}$. By hypothesis $J:=I_{c}(G)$ is linearly presented. Thus, by (c) in Proposition 3.4, there is a path $\mathcal{P}=\{\bar{C_{1}},\bar{C_{2}},\ldots,\bar{C_{k}}\}$ in $\mathcal{G}_{J}^{(C_{1},C_{2})}$ with $\bar{C_{1}}=C_{1}$, $\bar{C_{k}}=C_{2}$, and $\bar{C_{i}}\subset C_{1}\cup C_{2}$ for $i=1,\ldots k$. Furthermore, by Lemma 5.11(b), $G^{\prime}$ is a bipartite graph with a perfect matching and bipartition $(B_{1},B_{2})$ such that $|B_{1}|=|B_{2}|$, where $B_{1}:=C_{1}\setminus C_{2}$ and $B_{2}:=C_{2}\setminus C_{1}$. Since $G^{\prime}$ has a perfect matching, $G^{\prime}$ does not have isolated vertices. Also, by König’s theorem (Theorem 2.6), $G^{\prime}$ is a König graph. Hence, by Proposition 2.7(a), $G^{\prime}$ has a perfect matching $P$ of König type with the (P) property, since $G^{\prime}$ is unmixed. Then, $|V(G^{\prime})|=2|P|$ implies $|B_{1}|=|B_{2}|=|P|$, since $V(G^{\prime})=B_{1}\cup B_{2}$ and $B_{1}\cap B_{2}=\emptyset$. This implies, $|\bar{C_{i}}|=|C_{1}|=|A|+|P|$, since $G$ is unmixed, $C_{1}=A\cup B_{1}$ and $A\cap B_{1}=\emptyset$.
We claim that $A\not\subset\bar{C_{i}}$ for some $i$. Arguing by contradiction, we assume $A\subset\bar{C_{i}}$ for each $i\in\{1,\ldots,k\}$. Letting $\bar{C^{\prime}_{i}}:=\bar{C_{i}}\setminus A$, then $|\bar{C^{\prime}_{i}}|=|\bar{C_{i}}|-|A|=|P|$. We take $\epsilon\in P$. Since $\bar{C_{i}}$ is a vertex cover, we have that $\emptyset\neq\bar{C_{i}}\cap\epsilon=\bar{C^{\prime}_{i}}\cap\epsilon$, because $A\cap\epsilon\subset A\cap V(G^{\prime})=\emptyset$. Furthermore, $P$ is a matching, then
$|\bar{C^{\prime}_{i}}|\geq|\bar{C^{\prime}_{i}}\cap V(P)|=\sum_{\epsilon\in P}|\bar{C^{\prime}_{i}}\cap\epsilon|\geq|P|.$
But $|\bar{C^{\prime}_{i}}|=|P|$, then $1=|\bar{C^{\prime}_{i}}\cap\epsilon|=|\bar{C_{i}}\cap\epsilon|$ for each $\epsilon\in P$. On the other hand $G^{\prime}$ is not Cohen-Macaulay, then by Proposition 2.7(b), there is a $4$-cycle $(t_{1},t^{\prime}_{1},t_{2},t^{\prime}_{2},t_{1})$ such that $\epsilon_{1}=\{t_{1},t^{\prime}_{1}\},\epsilon_{2}=\{t_{2},t^{\prime}_{2}\}\in P$. Without loss of generality, we can assume $t_{1},t_{2}\in B_{1}$ and $t^{\prime}_{1},t^{\prime}_{2}\in B_{2}$, since $(B_{1},B_{2})$ is a bipartition of $G^{\prime}$. By a previous argument, $|\bar{C^{\prime}_{i}}\cap\epsilon_{1}|=1=|\bar{C^{\prime}_{i}}\cap\epsilon_{2}|$, since $\epsilon_{1},\epsilon_{2}\in P$. So, if $\bar{C^{\prime}_{i}}\cap\epsilon_{1}=\{t_{1}\}$, then $\bar{C^{\prime}_{i}}\cap\epsilon_{2}=\{t_{2}\}$ since $\{t^{\prime}_{1},t_{2}\}\in E(G^{\prime})$. Similarly, if $\bar{C^{\prime}_{i}}\cap\epsilon_{1}=\{t^{\prime}_{1}\}$, then $\bar{C^{\prime}_{i}}\cap\epsilon_{2}=\{t^{\prime}_{2}\}$ since $\{t_{1},t^{\prime}_{2}\}\in E(G^{\prime})$. This implies
$\bar{C_{i}}\cap(\epsilon_{1}\cup\epsilon_{2})=\{t_{1},t_{2}\}\mbox{ or }\bar{C_{i}}\cap(\epsilon_{1}\cup\epsilon_{2})=\{t^{\prime}_{1},t^{\prime}_{2}\},$
since $\bar{C_{i}}\cap\epsilon=\bar{C^{\prime}_{i}}\cap\epsilon$. In particular we have $C_{1}\cap(\epsilon_{1}\cup\epsilon_{2})=\{t_{1},t_{2}\}$ and $C_{2}\cap(\epsilon_{1}\cup\epsilon_{2})=\{t^{\prime}_{1},t^{\prime}_{2}\}$, since $t_{1},t_{2}\in B_{1}$ and $t^{\prime}_{1},t^{\prime}_{2}\in B_{2}$. Now, we let $j=\min\{i\mbox{ }|\mbox{ }\bar{C_{i}}\cap(\epsilon_{1}\cup\epsilon_{2})=\{t^{\prime}_{1},t^{\prime}_{2}\}\}$, then $j>1$ and $\{t^{\prime}_{1},t^{\prime}_{2}\}\ \subset\bar{C_{j}}\setminus\bar{C}_{j-1}$. This is a contradiction since $\{\bar{C_{j}},\bar{C}_{j-1}\}\in E(\mathcal{G}_{J}^{(C_{1},C_{2})})$. Therefore, there is $\bar{C}_{i^{\prime}}$ such that $A\not\subset\bar{C}_{i^{\prime}}$. Hence, to finish the proof, we can set $C:=\bar{C}_{i^{\prime}}$, since $A\not\subset\bar{C}_{i^{\prime}}\subset C_{1}\cup C_{2}$. ∎
Lemma 5.16.
Let $G$ be a graph without $3$- and $5$-cycles. If $G^{\prime}$ is an induced subgraph of $G$ with a perfect matching $P$ and $\overline{t}\in V(G)\setminus V(G^{\prime})$, then, the set
$B=\{t\in V(G)\mbox{ }|\mbox{ there is }t^{\prime}\in N_{G}(\overline{t})\mbox{such that }\{t,t^{\prime}\}\in P\}$
is a stable set.
Proof.
By contradiction assume there is $\{t_{1},t_{2}\}\in E(G)$ such that $t_{1},t_{2}\in B$. Then, $t_{1}\neq t_{2}$ and there are $t^{\prime}_{1},t^{\prime}_{2}\in N_{G}(\overline{t})$ such that $\{t_{1},t^{\prime}_{1}\},\{t_{2},t^{\prime}_{2}\}\in P$. If $t_{1}=t^{\prime}_{2}$, then $\{\overline{t},t_{1}\}\in E(G)$ and $(\overline{t},t^{\prime}_{1},t_{1},\overline{t})$ is $3$-cycle. A contradiction, since $G$ does not have $3$-cycle. Hence $t_{1}\neq t^{\prime}_{2}$. Consequently, $\{t_{1},t^{\prime}_{1}\}\neq\{t_{2},t^{\prime}_{2}\}$, since $t_{1}\neq t_{2}$ and $t_{1}\neq t^{\prime}_{2}$. Thus, $\{t_{1},t^{\prime}_{1}\}\cap\{t_{2},t^{\prime}_{2}\}=\emptyset$, since $P$ is a perfect matching. Then, $(\overline{t},t^{\prime}_{1},t_{1},t_{2},t^{\prime}_{2},\overline{t})$ is $5$-cycle. Thus, is a contradiction, since $G$ does not have $5$-cycles. Therefore, $B$ is a stable set. ∎
Theorem 5.17.
Let $G$ be an unmixed graph without $3$- and $5$-cycles. The following condition are equivalent.
  1. (a)
    $I_{c}(G)$ is linearly presented.
  2. (b)
    If $G^{\prime}=G[C_{1}\triangle C_{2}]$ is an unmixed graph, where $C_{1},C_{2}\in E(G^{\vee})$, $C_{1}\neq C_{2}$, then $G^{\prime}$ is a Cohen–Macaulay graph.
Proof.
(b) $\Rightarrow$ (a) This implication follows from Theorem 5.12.
(a) $\Rightarrow$ (b). We argue by contradiction, assume there are $C_{1},C_{2}\in E(G^{\vee})$ such that $C_{1}\neq C_{2}$, $G^{\prime}=G[C_{1}\triangle C_{2}]$ is unmixed and $G^{\prime}$ is not Cohen–Macaulay. Then, by Lemma 5.11(b), $G^{\prime}$ is a bipartite graph with a perfect matching $P$. By Theorem 5.15 , there is $\bar{C}\in E(G^{\vee})$ such that $A\not\subset\bar{C}\subset C_{1}\cup C_{2}$, where $A:=C_{1}\cap C_{2}$. So, there is $t_{1}\in A\setminus\bar{C}$, then $t_{1}\notin C_{1}\triangle C_{2}=V(G^{\prime})$ and $N_{G}(t_{1})\subset\bar{C}$ since $\bar{C}$ is a vertex cover. In particular $N_{G}(t_{1})\cap D=\emptyset$, where $D:=V(G)\setminus(C_{1}\cup C_{2})$, since $\bar{C}\subset C_{1}\cup C_{2}$. Furthermore, by Lemma 5.16
$B:=\{t\in V(G^{\prime})|\mbox{ there is }t^{\prime}\in N_{G}(t_{1})\mbox{ suchthat }\{t,t^{\prime}\}\in P\}$
is a stable set. Consequently, there is a maximal stable set $L^{\prime}$ of $G^{\prime}$ such that $B\subset L^{\prime}$. If $t\in L^{\prime}\subset C_{1}\triangle C_{2}$, then $t\notin C_{1}$ or $t\notin C_{2}$. Thus, $N_{G}(t)\subset C_{1}$ or $N_{G}(t)\subset C_{2}$, since $C_{1}$ and $C_{2}$ are vertex covers of $G$. So, $N_{G}(t)\cap D=\emptyset$. This implies, $L^{\prime}\cup D$ is a stable set. Also if $\{t_{1},\bar{t}\}\in E(G)$ with $\bar{t}\in L^{\prime}$, then there is $\{\bar{t},\bar{t^{\prime}}\}\in P$. Consequently, $\bar{t^{\prime}}\in B\subset L^{\prime}$. A contradiction, since $\bar{t},\bar{t^{\prime}}\in L^{\prime}$ and $L^{\prime}$ is a stable set. Hence, $L=L^{\prime}\cup D\cup\{t_{1}\}$ is a stable set, since $N_{G}(t_{1})\cap D=\emptyset$. Furthermore, $L^{\prime}\cap D=\emptyset$ and $t_{1}\notin L^{\prime}\cup D$, since $L^{\prime}\subset C_{1}\triangle C_{2}$ and $t_{1}\in A$. Thus, $|L|=|L^{\prime}|+|C|+1$. On the other hand, $L_{1}:=V(G)\setminus C_{1}$ is a maximal stable set, since $C_{1}\in E(G^{\vee})$. Notice $L_{1}=B_{2}\cup D$ where $B_{2}:=C_{2}\setminus C_{1}$, since $V(G)=(C_{1}\cup C_{2})\cup D$. In particular $B_{2}$ is a stable set and $B_{2}\subset V(G^{\prime})$. Then, $|L^{\prime}|\geq|B_{2}|$ implies
$|L|=|L^{\prime}|+|D|+1\geq|B_{2}|+|D|+1\geq|L_{1}|+1,$
This is a contradiction, since $G$ is unmixed. ∎
Proposition 5.18.
Let $U,U_{1},U_{2},U_{3},U_{4},U_{5}$ be the families of graphs defined in Section 1. Then,
$U_{5}\subset U_{4}\subset U_{3}\subset U_{2}\subset U_{1}\subset U$
and all inclusions are strict.
Proof.
By Theorem 5.13, Proposition 3.4, Theorem 5.4 and Lemma 5.3 all inclusions hold. To show that all inclusions are strict note that Example  6.5 is in $U_{4}\setminus U_{5}$, Example  6.7 is in $U_{3}\setminus U_{4}$ and Example  6.6 is in $U_{2}\setminus U_{3}$. Now we take a homogeneous $k$-partite graph $G$ with $k\geq 3$ such that $G$ is not a complete graph. Then, by (b) and (c) in Proposition 3.14, $G\in U_{1}$. Furthermore, as $G$ is not a complete graph, one has $G=G_{1}*\cdots*G_{k}$, with $|G_{i}|\geq 2$. Thus, $G\notin U_{2}$, since the vertices of $G_{1}$ are duplicated. Hence, $G\in U_{1}\setminus U_{2}$. Finally if $G^{\prime}$ is a homogeneous complete bipartite graph with $|V(G)|\geq 4$, then $G^{\prime}$ has a $4$-cycle. Therefore, by (a) and (c) in Proposition 3.14, $G^{\prime}\in U\setminus U_{1}$. ∎
6. Examples
Example 6.1.
Let $G=C^{5}$ be a $5$-cycle with vertex set $V(G)=\{t_{1},\ldots,t_{5}\}$ and let $C$ be the minimal vertex cover $\{t_{1},t_{3},t_{4}\}$ of $G$. Then,
$\ t_{2}\notin C,\ t_{3}\in C,\ \{t_{2},t_{3}\}\in E(G),\ N_{G}(t_{3})\setminus C=\{t_{2}\},\ \mbox{and}$
$N_{G}(t_{3})=\{t_{2},t_{4}\}\not\subset N_{G}[t_{2}]=\{t_{1},t_{2},t_{3}\}.$
More generally one has $N_{G}(t_{i})\not\subset N_{G}[t_{j}]$ for any $\{t_{i},t_{j}\}\in E(G)$. Note that all minimal vertex covers of $G$ satisfy the exchange property. Thus, the converse of Proposition 3.6 does not hold. If $C^{s}$ is a cycle of length $s$, then ${\rm v}(I_{c}(C^{s}))=\lfloor\frac{s}{2}\rfloor$ [26, Proposition 3.3]. In our case ${\rm v}(I_{c}(G))=2$. The graph $\mathcal{G}_{I_{c}(G)}$ of $I_{c}(G)$ is again a $5$-cycle and, by Proposition 4.10, the columns of the linear syzygy matrix of $I_{c}(G)$ are linearly independent.
Example 6.2.
Let $G=C^{7}$ be a $7$-cycle with vertex set $V(G)=\{t_{1},\ldots,t_{7}\}$ and let $S=\mathbb{Q}[V(G)]$ be a polynomial ring over the field $\mathbb{Q}$. The ideal $I_{c}(G)$ is minimally generated by monomials of degree $4$. By Theorem 5.13, $I_{c}(G)$ is linearly presented. But using Macaulay2222 [14], $I_{c}(G)$ does not have a linear resolution. Furthermore, by Theorem 3.5, ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1=3$.
Example 6.3.
Let $G_{1}$ and $G_{2}$ be two vertex disjoint 5-cycles. The join $G=G_{1}*G_{2}$ consists of $G_{1}\cup G_{2}$ and all the lines joining $G_{1}$ with $G_{2}$. The complement $\overline{G}$ of $G$ is $G_{1}\cup G_{2}$ and $G$ is a codi-graph but $G$ is not a cograph because $G_{1}$ is an induced connected subgraph of $G$ whose complement is connected. By Example 6.1 and Proposition 3.12, one has ${\rm v}(I_{c}(G))=2+5=7$.
Example 6.4.
Let $C^{3}$ and $C^{5}$ be two vertex disjoint odd cycles of length $3$ and $5$, respectively, and let $G=C^{3}*C^{5}$ be their join. Then, ${\rm v}(I_{c}(C^{3}))+|V(C^{5})|=1+5=6$ and ${\rm v}(I_{c}(C^{5}))+|V(C^{3})|=2+3=5$. Thus, by Proposition 3.12, one has ${\rm v}(I_{c}(G))=5$. As $G$ is a codi-graph, one has ${\rm reg}(S/I_{c}(G))=|V(G)|-2=6$ (Proposition 3.9).
Example 6.5.
Let $G$ be a graph consisting of an induced $4$-cycle $C^{4}=\{t_{1},t_{2},t_{3},t_{4},t_{1}\}$, and edges $\epsilon_{1}=\{t_{1},t^{\prime}_{1}\}$, $\epsilon_{2}=\{t_{2},t^{\prime}_{2}\}$, $\epsilon_{3}=\{t_{3},t^{\prime}_{3}\}$ and $\epsilon_{4}=\{t_{4},t^{\prime}_{4}\}$. Then, $P=\{\epsilon_{1},\epsilon_{2},\epsilon_{3},\epsilon_{4}\}$ is a perfect matching with the (P) property. Consequently, by Proposition 5.1, $\alpha_{0}(G)=4$. Thus, $P$ is a perfect matching of König type and $G$ is König. Therefore, by Proposition 2.7(b) and Theorem 5.8, $G$ is Cohen–Macaulay and $I_{c}(G)$ is linearly presented.
Example 6.6.
Let $G$ be the graph depicted in Figure 1 consisting of: a complete subgraph $\mathcal{K}_{4}$ with $V(\mathcal{K}_{4})=\{t_{1},t_{2},t_{3},t_{4}\}$ and four $4$-cycles $C^{(1)}=\{t_{1},t_{2},t_{6},t_{5},t_{1}\}$, $C^{(2)}=\{t_{1},t_{8},t_{7},t_{3},t_{1}\}$, $C^{(3)}=\{t_{2},t_{9},t_{10},t_{4},t_{2}\}$ and $C^{(4)}=\{t_{3},t_{4},t_{8},t_{9},t_{3}\}$. Using Macaulay2222 [14], one obtains that $G$ is an unmixed graph with $\alpha_{0}(G)=6$ whose minimal vertex covers are:
$C_{1}=\{t_{2},t_{3},t_{4},t_{5},t_{8},t_{9}\},\ C_{2}=\{t_{2},t_{3},t_{4},t_{5},t_{8},t_{10}\},\ C_{3}=\{t_{1},t_{2},t_{3},t_{5},t_{8},t_{10}\},$
$C_{4}=\{t_{1},t_{2},t_{3},t_{6},t_{8},t_{10}\},\ C_{5}=\{t_{1},t_{3},t_{4},t_{6},t_{8},t_{9}\},\ C_{6}=\{t_{1},t_{3},t_{4},t_{6},t_{7},t_{9}\},$
$C_{7}=\{t_{1},t_{2},t_{4},t_{6},t_{7},t_{9}\}\mbox{ and }C_{8}=\{t_{1},t_{2},t_{4},t_{5},t_{7},t_{9}\}.$
Hence, $\mathcal{G}_{I_{c}(G)}$ consists of two disjoint path ${\mathcal{P}}_{1}=\{C_{1},C_{2},C_{3},C_{4}\}$ and ${\mathcal{P}}_{2}=\{C_{5},C_{6},C_{7},C_{8}\}$. Therefore, $\mathcal{G}_{I_{c}(G)}$ is not connected. Also, the exchange edges of $G$ are:
$\{t_{9},t_{10}\},\ \{t_{1},t_{4}\},\ \{t_{5},t_{6}\},\ \{t_{7},t_{8}\}\mbox{ and }\{t_{2},t_{3}\}.$
On the other hand note that the graph $G$ has no duplicated vertices. Then, by Lemma 5.3, every $4$-cycle has an edge without the (P) property. Furthermore, the regularity of $S/I_{c}(G)$ is $6$ and ${\rm v}(I_{c}(G))=\alpha_{0}(G)-1=5$.
$t_{3}$$t_{1}$$t_{5}$$t_{4}$$t_{2}$$t_{6}$$t_{8}$$t_{7}$$t_{9}$$t_{10}$
Figure 1. Unmixed graph without duplicated vertices, not König and $\mathcal{G}_{I_{c}(G)}$ not connected.
Example 6.7.
Let $G$ be the graph depicted in Figure 2 consisting of two cycles $C^{3}=\{t_{1},t_{2},t_{3},t_{1}\}$ and $C^{4}=\{t_{1},t_{2},t_{7},t_{6},t_{1}\}$ and two edges $\{t_{4},t_{5}\}$ and $\{t_{3},t_{4}\}$. Using Macaulay2222 [14], one obtains that $G$ is an unmixed graph with $\alpha_{0}(G)=4$ whose minimal vertex covers are:
$C_{1}=\{t_{2},t_{3},t_{6},t_{5}\},\ C_{2}=\{t_{2},t_{3},t_{6},t_{4}\},\ C_{3}=\{t_{1},t_{2},t_{4},t_{6}\},\ C_{4}=\{t_{1},t_{2},t_{4},t_{7}\},$
$C_{5}=\{t_{1},t_{3},t_{7},t_{4}\}\mbox{ and }C_{6}=\{t_{1},t_{3},t_{7},t_{5}\}.$
The graph $G$ has duplicated exchange edges because the edge $\{t_{4},t_{5}\}$ of $G$ is an exchange edge appearing in $(C_{1}\setminus\{t_{5}\})\cup\{t_{4}\}=C_{2}$ and $(C_{5}\setminus\{t_{4}\})\cup\{t_{5}\}=C_{6}$. Letting $J=I_{c}(G)$, then $\mathcal{G}_{J}$ is the path $\{C_{1},C_{2},C_{3},C_{4},C_{5},C_{6}\}$ with $6$ vertices. Hence, $\mathcal{G}_{J}$ is connected. Furthermore, the subgraph $\mathcal{G}_{J}^{(C_{1},C_{6})}$ of $\mathcal{G}_{J}$ consist of two isolated vertices $C_{1}$ and $C_{6}$, since $t_{4}\notin C_{1}\cup C_{6}$ and $t_{4}\in\cap_{i=2}^{5}C_{i}$. In particular $\mathcal{G}^{(C_{1},C_{6})}$ is not connected. Therefore, by Proposition 3.4, $I_{c}(G)$ is not linearly presented. On the other hand, $\mathcal{G}_{J}^{(C_{2},C_{5})}$ is the path $\{C_{2},C_{3},C_{4},C_{5}\}$ and the exchange edges of this path do not form a matching of $G$. This means that Proposition 4.2(b) does not extend to paths of length $3$.
$t_{3}$$t_{1}$$t_{2}$$t_{6}$$t_{7}$$t_{4}$$t_{5}$
Figure 2. Unmixed graph with $\mathcal{G}_{J}$ connected and $\mathcal{G}_{J}^{(C_{1},C_{6})}$ not connected.
Example 6.8.
Let $G$ be the graph of Figure 2 and let $H$ be the subgraph of $G$ induced by the vertices $t_{1},t_{2},t_{3},t_{6},t_{7}$. Then, $H$ consists of two cycles: $\{t_{1},t_{2},t_{3},t_{1}\}$ and $\{t_{1},t_{2},t_{7},t_{6},t_{1}\}$. Using Macaulay2222 [14], one obtains that $G$ is an unmixed graph with $\alpha_{0}(G)=3$ whose minimal vertex covers are:
$C_{1}=\{t_{1},t_{3},t_{7}\},\ C_{2}=\{t_{1},t_{2},t_{7}\},\ C_{3}=\{t_{1},t_{2},t_{6}\}\mbox{ and }C_{4}=\{t_{2},t_{3},t_{6}\}.$
Letting $J=I_{c}(H)$, the graph $\mathcal{G}_{J}$ of $J$ is the path (with $4$ vertices) $\mathcal{P}=\{C_{1},C_{2},C_{3},C_{4}\}$. Furthermore, the subgraph $\mathcal{G}_{J}^{(C_{i},C_{j})}$ of $\mathcal{G}_{J}$ is connected, for all $1\leq i<j\leq 4$. Hence, by Proposition 3.4, $J$ is linearly presented. But $H^{\prime}:=H[C_{1}\triangle C_{4}]=\{t_{1},t_{2},t_{7},t_{6},t_{1}\}$ is a $4$-cycle, then $H^{\prime}$ is an unmixed König graph and it is not Cohen–Macaulay. Therefore, the converse of Theorem 5.12 is not true.
Appendix A Procedures
Procedure A.1.
Computing the graph $\mathcal{G}_{I_{c}(G)}$ of the ideal of covers $I_{c}(G)$ of a graph $G$ using Macaulay2222 [14]. This procedure corresponds to Example 6.6. One can compute other examples by changing the polynomial ring $S$ and the generators of the ideal $I$.
restart
load "EdgeIdeals.m2"
--The ti’s correspond to the vertices of the graph G
--and the yi’s correspond to the vertex covers of G
--These should be changed to run other examples
S=QQ[t1,t2,t3,t4,t5,t6,t7,t8,t9,t10,y1,y2,y3,y4,y5,y6,y7,y8,y9,y10]
--This is the edge ideal of G
I=monomialIdeal(t1*t2,t2*t3,t3*t4,t1*t4,t1*t3,t2*t4,t1*t5,t5*t6,t6*t2,
t1*t8,t2*t9,t3*t7,t4*t10,t7*t8,t10*t9,t8*t9,t3*t9,t4*t8)
--This is the ideal of covers of G
Ic=dual I
--This checks whether or not G is unmixed
degrees Ic
M=coker gens gb Ic
pdim M
--The syzygy matrix of Ic(G) is:
syz gens Ic
--G has 8 minimal vertex covers corresponding to y1,...,y8
f=matrix{{y1},{y2},{y3},{y4},{y5},{y6},{y7},{y8}}
J=ideal(transpose(f)*(syz gens Ic))
--Eliminate the generators of J of degree > 2 to obtain:
J1=ideal(apply(flatten entries gens J,x-> if degree(x)=={2}
 then x else 0))
--number of independent linear syzygies of Ic(G):
numgensJ1=#(set(apply(flatten entries gens J,x->
if degree(x)=={2} then x else 0))-{0})
fsub=gens sub(ideal(toList (set(apply(flatten entries gens J,
x-> if degree(x)=={2} then x else 0))-{0})),
{t1=>1,t2=>1,t3=>1,t4=>1,t5=>1,t6=>1,t7=>1,t8=>1,t9=>1,t10=>1})
--This gives the graph of Ic(G) if the columns of the
--linear syzygy matrix of Ic(G) are linearly independent:
fsub1=monomialIdeal gens gb ideal apply(#(flatten entries fsub),
n->someTerms((flatten entries fsub)#n,1,1)*
someTerms((flatten entries fsub)#n,0,1))
GIc=graph(fsub1)
isBipartite GIc
allOddHoles(GIc)
isConnected(GIc)
L=ass Ic
p=(n)->gens gb ideal(flatten mingens(quotient(Ic,L#n)/Ic))
minA=monomialIdeal(apply(0..#L-1,p))
--This computes the v-number of Ic(G)
vnumber=min flatten degrees minA
--The regularity of Ic(G)
regularity M
--minimal free resolution of Ic(G)
Mres=res M
--This is the syzygy matrix of Ic(G)
Mres.dd_2
--Linearly independent linear syzygies of Ic(G)
submatrix(Mres.dd_2,{0..5})
--rank of the linear syzygy matrix
--this rank is number of minimal vertex covers -1
--if Ic(G) is linearly presented
rank submatrix(Mres.dd_2,{0..5})
--Minors of the linear syzygy matrix of Ic(G) if
--columns are linearly independent
minors(6,submatrix(Mres.dd_2,{0..5}))
--codim >=2 if Ic(G) is linearly presented:
codim minors(6,submatrix(Mres.dd_2,{0..5}))
--Now we use the edge ideal I to compute the graph of Ic(G)
--using a method that works in general
G=flatten apply(ass(I),x-> entries gens gb (x))
A1=toList subsets(G,2)
B1=toList set apply(A1,x->toList x)
--This gives the edges of the graph of Ic(G)
graphIc=apply(0..#B1-1,n-> if #(set((B1#n)#0)*set((B1#n)#1))==
codim(I)-1 then  (set((B1#n)#0),set((B1#n)#1)) else set{})
graph1=toList set graphIc
--This is the number of edges of the graph of Ic(G)
#(set graph1)-1
--If the following two numbers are equal
--the linear syzygy matrix of Ic(G) has linearly independent columns
#(set graph1)-1==numgensJ1
Acknowledgments.
Macaulay2222 [14] was used to implement algorithms to compute the graph of ideals of covers and the v-number of monomial ideals.